### v1

# Unlocking Open-Set Language Accessibility in Vision Models

Abstract
Visual classifiers offer high-dimensional feature representations that are challenging to interpret and analyze. Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior. We propose a simple, yet powerful method for reformulating any visual classifier so that it can be accessed with open-set text queries without compromising its original performance. Our approach is label-free, efficient, and preserves the underlying classifier’s distribution and reasoning processes. We thus unlock several text-based interpretability applications for any classifier. We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language. In both applications, we achieve state-of-the-art results, greatly outperforming existing works. Our method enables text approaches for interpreting visual classifiers.
Machine Learning, ICML
1 Introduction
Visual classifiers inherently and purely provide visual signals, which are high-dimensional dense vector representations of images that are hard to interpret and dissect, limiting attribution maps as the main medium for interpretation. Attribution maps provide general and high-level information, indicating where in the image the most important region is for the prediction. Nevertheless, text is significantly more expressive, human-friendly, specific, and unlocks a variety of other interpretability applications that require language (Tewel et al., 2021; Koh et al., 2020; Oikarinen & Weng, 2023; Song et al., 2022; Minderer et al., 2022). Therefore, a question that arises is how can we talk or access those visual classifiers via open-set text? Performing this allows us to have a text interface that can be used to query visual classifiers. So far, this has only been achieved with generic vision-language models such as CLIP-based ones (Radford et al., 2021), which learn a shared vision-language space and thus enables any text to query the visual model directly. However, the deep learning field extends far beyond CLIP. For example, industries—being the primary end-users of such models—often rely on custom visual models tailored to their specific needs. This creates an application bottleneck: any task involving both visual and textual interpretations becomes restricted to CLIP or similar approaches, limiting flexibility, forcing reliance on such models, and inheriting their limitations.
Figure 1: We formulate any classifier to allow querying it and measuring its association with any text, while maintaining its performance and distribution. $f$ is the feature vector of an image.
In this work, we lift this restriction by proposing a method to reformulate any visual classifier such that it can be accessed via open-set text (Figure 1). Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, it is faithful to the classifier and preserves its underlying distribution and reasoning process. Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. In ImageNet-1K (Deng et al., 2009), there are a total of $1,000$ labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information. Specifically, these embeddings also semantically encode words that are close to the class name within the text embedding space. For the “goldfish” example, the closest words could be “freshwater”, “fins” and “orange”. Our method learns to map images into this text embedding space, thus associating both the class name and its surrounding semantic space with the image. This is accomplished through a trainable MLP that projects the visual features into the text embedding space, and is trained by aligning its output distribution across all classes with the original classifier’s output distribution, while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with any text query beyond the class names.
We demonstrate the effectiveness of our method with two primary applications. 1) We build label-free concept bottleneck models for any classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the new linear classifier remains faithful to the original classifier. 2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
In summary, our contributions are as follows:
  • We propose an efficient, label-free method to reconfigure any visual classifier such that it can be queried with open-set text queries, while also being faithful to the classifier.
  • We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free, faithful zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language.
  • Our method sets new state-of-the-art results outperforming existing works, including those which use CLIP as the visual encoder, although being trained with 400$\times$ less images and 400,000$\times$ less captions, thus efficiently learning faithful text interfaces to interpret any visual classifier.
2 Related Work
While CLIP is the predominant approach to interpret vision models through language (Oikarinen & Weng, 2023), there exist works that try to decode visual features of models beyond CLIP. DeVIL (Dani et al., 2023) trains an autoregressive text generator to map visual features from different layers of a classifier into image captions, leveraging annotated image-caption pairs as ground-truth data. Similarly, Natural Language Explanations (NLEs) (Park et al., 2018; Kayser et al., 2021; Sammani et al., 2022) use annotated textual explanations in place of conventional captions. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, and are therefore not faithful to the classifier. ZS-A2T (Salewski et al., 2023) maps attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective), and thus cannot be applied to visual classifiers. Our work on the other hand decodes visual features of any classifier in a zero-shot manner without requiring any annotated data, while also maintaining faithfulness to the original classifier.
Text-to-Concept (T2C) (Moayeri et al., 2023) is the closest related work to ours, where a linear layer is trained to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Since the linear layer maps features to the CLIP space, T2C is strongly biased towards interpreting the CLIP model rather than the original classifier. As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image. Moreover, T2C relies on CLIP supervision. In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution, ensuring faithfulness to it.
3 Method
Figure 2: An overview of our method for reformulating visual classifiers. (a) The process of training the MLP mapping between vision and text space. (b) The process of inference with adapted visual classifier. The text encoder acts as a linear classifier weight generator in our re-formulated classification process. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
Consider an image $I$ and a visual classifier $F$ composed of a visual feature extractor $F_{v}$ and a linear classifier $W$. Note that $F$ can be of any architecture. $F_{v}$ embeds $I$ into an $n-$dimensional feature vector $f\in\mathbb{R}^{n}$. That is, $f=F_{v}(I)$. The linear classifier $W\in\mathbb{R}^{n\times K}$ takes $f$ as input and outputs a probability distribution $o$ for the image across $K$ classes. That is, $o=\text{softmax}(f.W)\in\mathbb{R}^{K}$. For ImageNet-1K, $K=1000$. Consider also any off-the-shelf text encoder $T$ which takes in an input text $l$ and embeds it into a $m-$dimensional vector representation $u\in\mathbb{R}^{m}$. That is, $u=T(l)$. Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
We propose to learn a simple light-weight MLP mapping function which projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$. Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient. We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}” where {class} represents the class name in text format. This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$. Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$. We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$. Equivalently, this can be performed as a single matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class.
The most straightforward approach to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions). Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original distribution $o$ through the cross-entropy loss. For a single sample, the loss is given by:
$L=-\sum_{i=1}^{K}o_{i}\log\left(\frac{e^{s_{i}}}{\sum_{j=1}^{K}e^{s_{j}}}\right)\;.$ (1)
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification. This makes our approach label-free and preserves the distribution and reasoning process of the classifier $F$. We provide a PyTorch-like pseudocode of our training approach in Listing LABEL:lst:pseudo, and illustrate it in Figure 2(a). It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions or hierarchies (see Section F in the Appendix for more information).
After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features. In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the newly formulated classification task, defined as argmax($\tilde{f}.U^{T}$). This is shown in Figure 2(b). We show later that this weight generating process is important to build faithful concept bottleneck models.
# text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
# classifier: linear classifier weights of a frozen vision_encoder, shape (visual_dim, num_classes)
# mlp: trainable MLP from visual_dim -> text_dim
# images: batch of B images, shape (N, 3, height, width)
visual_feats = vision_encoder(images) # (N, visual_dim)
logits = visual_feats @ classifier # (N, num_classes)
original_dist = softmax(logits, dim=-1) # (N, num_classes)
mapped_feats = mlp(visual_feats) # (N, text_dim)
mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
text_feats = l2_norm(text_feats) # (N, text_dim)
pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
# cross entropy with original model’s soft distribution
loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
loss.backward() # only mapper parameters are updated
Listing 1: PyTorch-like pseudocode for our method
4 Applications
In this section, we show how our method can be used to build text-based interpretability applications. At application, all model components (including the MLP) are frozen.
4.1 Zero-Shot Concept Bottleneck Models
Concept Bottleneck Models (CBMs) (Koh et al., 2020) are a class of inherently interpretable models and have recently attracted significant attention. They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions. In (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score which represents the concept activation to the image. In (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of dense features. Recently, the Label-Free CBMs (LF-CBMs) family uses CLIP to perform the concept discovery step without annotated image-concept data, by either using CLIP as a model to provide ground-truth image-concept similarities to train a layer which performs concept discovery (Oikarinen et al., 2023), or by querying the image features from a set of predefined concepts within the CLIP space (Yang et al., 2022; Panousis et al., 2023) and using the cosine similarity scores between the concepts and the image as concept activations. However, there are major issues in the current literature of CBMs. Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model. Secondly, all CBM approaches to date require training a linear probe on the concept activations which 1) hinders an on-the-fly ready-to-deploy CBMs, and 2) disconnects $W^{con}$ from the original feature-based classifier $U$, since $W^{con}$ is trained from scratch in a different space using ground-truth data and is hence not faithful to the original model, and 3) requires training a new $W^{con}$ classifier for every different set of concepts, hindering flexibility to the choice of concepts.
Our method allows us to solve all the aforementioned issues and formulate zero-shot CBMs for any classifier. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ represents the classification weights of the newly formulated classifier (the output of the text encoder $T$ for the class prompts). We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works (Rao et al., 2024), and without loss of generality, we use the $Z=20K$ most common words in English [^2] https://github.com/first20hours/google-10000-english as our concept set. These are general concepts that represent world knowledge and are not tailored towards any specific dataset. We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$. That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$. By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$. For an image $I$, we extract its visual features $f$ and use MLP to map them to $\tilde{f}$ which now lies in the text embedding space. That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$. Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to. That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and $C$. The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts. We provide an illustration in Figure 3.
Figure 3: Our proposed approach to concept discovery in CBMs.
We now show how to build the classifier $W^{con}$ that takes in the concept activations and outputs a distribution $S_{cn}$ over classes. We build $W^{con}$ in a zero-shot manner. Here, zero-shot indicates that no training is required to map concept activations to classes. Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space. Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name. Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the linear classifier $U$ to obtain the new weights for $W^{con}$. That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$. Therefore, the weights of $W^{con}$ represent how similar the class name is to all the concepts. In total, the output distribution $S_{cn}$ of the CBM is defined as:
[…]
By observing Eq. 2, we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, all we have to do is plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs are faithful to the original classifier. By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner, for any classifier, while also maintaining faithfulness to the original classifier and being flexible in the concept set.
4.2 Zero-Shot Decoding of Visual Features into Text
In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into natural language sentence. This offers a text-based interpretation of what the visual features contain. We adapt the method introduced in (Tewel et al., 2021) for our purpose. Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-self pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li & Liang, 2021) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual” tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. An overview of this process is shown in Figure 4. As this process is not the core contribution of our work, we leave more details to Section E of the appendix, which also includes a detailed illustration in Figure 9.
Figure 4: To decode visual features into natural language using any pretrained language decoder (e.g., GPT-2), we apply prefix tuning while keeping the language decoder fixed, generating text that maximizes the similarity with the visual features. The figure is shown for iteration $j=1$
5 Experiments
We first provide results on our newly formulated visual classifiers. We use the ImageNet-1K dataset due to the widespread availability of visual classifiers trained and evaluated on it. We apply our method on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He et al., 2015), Wide ResNets (Zagoruyko & Komodakis, 2016), ResNeXts (Xie et al., 2016), ShuffleNetv2 (Ma et al., 2018), EfficientNetv2 (Tan & Le, 2021), Densely Connected Networks (DenseNets) (Huang et al., 2016), ConvNeXts (Liu et al., 2022) and ConvNeXtv2 (Woo et al., 2023). For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (Dosovitskiy et al., 2021), DINOv2 (Oquab et al., 2024), BeiT (Bao et al., 2022), the hybrid Convolution-Vision Transformer CvT (Wu et al., 2021), Swin Transformer (Liu et al., 2021b) and Swin Transformer v2 (Liu et al., 2021a). All models are pretrained on ImageNet-1K from the PyTorch [^3] https://pytorch.org/vision/stable/models.html and HuggingFace (Wolf et al., 2020) libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe (Vryniotis, 2021). Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. Both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder (Wang et al., 2020), with ablation studies on alternative text encoders shown in Section B of the Appendix. For implementation details, we refer to Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We refer to Appendix Section C for results on other models.
Model Top-1 Orig. $\Delta$
ResNet50 75.80 76.13 --0.33
[…]
ResNet101 77.19 77.37 --0.18
WideResnet50 78.35 78.47 --0.12
[…]
DenseNet161 77.04 77.14 --0.10
DenseNet169 75.46 75.60 --0.14
EfficientNetv2-S 84.04 84.23 --0.19
EfficientNetv2-M 84.95 85.11 --0.16
[…]
ConvNeXt-Small 83.42 83.62 --0.20
ConvNeXt-Base 83.88 84.06 --0.18
[…]
ResNeXt50-32x4d 77.44 77.62 --0.18
[…]
ResNeXt101-64x4d 83.13 83.25 --0.12
ResNeXt101-32x8d 79.10 79.31 --0.21
ViT-B/16 80.70 81.07 --0.37
[…]
ViT-L/32 76.72 76.97 --0.25
ViT-L/16 79.56 79.66 --0.10
Swin-Base 83.22 83.58 --0.36
Swinv2-Base 83.72 84.11 --0.39
BeiT-B/16 84.54 85.06 --0.52
BeiT-L/16 87.22 87.34 --0.12
DINOv2-B 84.40 84.22 +0.18
ConvNeXtV2-B 84.56 84.73 --0.17
[…]
Table 1: Performance comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents the difference between them $\Delta$ = Top-1 - Orig.
Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2. As baselines, we compare with LF-CBMs (Oikarinen et al., 2023), LaBo (Yang et al., 2022), CDM (Panousis et al., 2023) and the recent state-of-the-art DN-CBMs (Rao et al., 2024). All these methods are supervised and use CLIP-based models. Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art. Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples. The best results are obtained by the ConvNeXtv2 and ViT-L/16 models, which achieve a top-1 accuracy of $86.3$. All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
Method Model Top-1
Supervised CBMs
LF-CBM CLIP ResNet50 67.5
LF-CBM CLIP ViT-B/16 75.4
LaBo CLIP ResNet50 68.9
LaBo CLIP ViT-B/16 78.9
CDM CLIP ResNet50 72.2
CDM CLIP ViT-B/16 79.3
DCLIP CLIP ViT-B/16 68.0
DN-CBM CLIP ResNet50 72.9
DN-CBM CLIP ViT-B/16 79.5
Zero-Shot CBMs (Ours)
ZS-CBM ResNet50 73.9
[…]
ZS-CBM ResNet101 75.3
[…]
ZS-CBM WideResNet50 76.9
[…]
ZS-CBM DenseNet121 69.9
ZS-CBM DenseNet161 75.2
ZS-CBM EfficientNetv2-S 83.0
ZS-CBM EfficientNetv2-M 83.9
ZS-CBM ConvNeXt-Small 81.9
ZS-CBM ConvNeXt-Base 82.8
ZS-CBM ViT-B/32 73.3
ZS-CBM ViT-B/16 79.3
[…]
ZS-CBM Swin-Base 82.2
ZS-CBM Swinv2-Base 82.6
[…]
ZS-CBM BeiT-B/16 83.0
ZS-CBM DINOv2-B 82.6
[…]
ZS-CBM BeiT-L/16 86.2
[…]
Table 2: Performance of Zero-Shot CBMs (ZS-CBM) on ImageNet validation set for several classifiers.
Figure 5: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction. All examples use the 20K concept set, except for the first which uses the ImageNet LF-CBM concept set
Figure 6: Global class-wise interpretability analysis with our Concept Bottleneck Model. We highlight in yellow the top concepts in ”hammerhead shark” that are not present in ”tiger shark”, and therefore distinctive to ”hammerhead shark”.
In Figure 5, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. By observing the second example, we see that the image is predicted as a “goose” because it has duck-like features, it is in the size of a swan, and it has bird and pigeon-like features. Similarly, we can see that the third image is a “spotted salamander” due to its lizard-like structure, a head that looks like the frog’s head, a snake-like tail and spots like a leopard. In the last image, we can see that the prediction is “ostrich” due to its camel-like hump, mammal-like size, piegeon-like body, the horse-riding-like feature, and is a type of a bird. In Figure 6, we present a probability distribution of global class-wise concepts. These are concepts detected for all images of a specific class, along with their frequency. We consider two semantically similar classes but distinctively different: “hammerhead shark” and a “tiger shark”. We highlight in yellow the top concepts in “hammerhead shark” that are not present in “tiger shark”. These concepts are “harpoon” and “lobster hammer”, both which are distinctive to the head of the hammerhead shark and drive its prediction.
Figure 7: Qualitative examples of decoded visual features from different visual classifiers, into natural text.
Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features. Since the ground-truth content of visual features is unknown, we rely on COCO captions (Lin et al., 2014) for evaluation. While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose. Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person). Therefore, it also serves as a way to evaluate generalization of our method to other datasets, given that we only used the ImageNet class names for training. Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark. We present results on the widely used “Karpathy test split” with various vision classifiers.
As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap (Tewel et al., 2021) and ConZIC (Zeng et al., 2023), both which use CLIP. For evaluation, we employ standard natural language generation (NLG) metrics: BLEU-4 (B@4) (Papineni et al., 2002), METEOR (M) (Banerjee & Lavie, 2005), ROUGE-L (R-L) (Lin, 2004), CIDEr (C) (Vedantam et al., 2014), and SPICE (S) (Anderson et al., 2016). Results are shown Table 3. ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems. Even with a simple ResNet-50 vision encoder trained on ImageNet-1K (1.2 million images), our approach outperforms the baseline methods on CIDEr and SPICE, despite the latter utilizing the significantly more powerful CLIP vision encoder, trained on 400 million image-text pairs. We also present qualitative examples of the decoded visual features for different models in Figure 7. We can see from the first example, that BeiT-L/16 contains both the vegetables and dog in its features, while ConvNexTv2 only focuses on the vegetables, and ViT-B/16 and DINOv2 only focus on on the dog. Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason. This application allows a user-friendly way of interpreting visual features, applicable even to a layman user.
Model B4 M R-L C S
ZeroCap 2.6 11.5 14.6 5.5
ConZIC 1.3 11.5 12.8 5.2
DenseNet161 1.50 10.2 20.4 15.8 6.3
ResNet50 1.43 10.2 20.3 15.9 6.2
WideResNet50 1.40 10.2 20.4 16.0 6.4
[…]
DINOv2-Base 1.50 10.7 21.0 17.3 6.7
[…]
BeiT-L/16 1.50 10.6 20.9 17.6 6.9
[…]
Table 3: Zero-Shot Image Captioning Performance
Model B4 M R-L C S
ZeroCap 2.6 11.5 14.6 5.5
ConZIC 1.3 11.5 12.8 5.2
Ours
DenseNet161 4.20 12.5 30.1 17.0 6.6
ResNet50 4.10 12.5 30.1 17.0 6.6
[…]
ResNet101v2 4.30 12.6 30.1 18.0 6.8
ConvNeXt-Base 4.50 12.7 30.3 18.1 6.9
[…]
EfficientNetv2-S 4.40 12.7 30.4 18.6 6.9
[…]
BeiT-B/16 4.50 12.8 30.3 18.9 7.1
DINOv2-Base 4.60 13.0 30.7 18.7 7.1
Table 4: Composition Captioning Performance across models.
Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this by revisiting an old image captioning paradigm termed as compositional captioning (Kulkarni et al., 2011), and later revived with deep learning methods (Lu et al., 2018). In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3, and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 4.1. We also add a list of the most common verbs [^4] https://github.com/datmt/English-Words-Updated in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a very few in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini (OpenAI, 2024) as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances. We provide details on the exact prompt we used in Appendix Section G. Results are shown in Table 4. While results on CiDEr and SPICE are incremental compared to results in Table 3, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3 compared to baseline methods. Also note that since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating sentences in a specific style (e.g., humorous, positive, or negative).
6 Conclusion and Future Directions
In this work, we presented a method to transform visual classifiers such that they can be queried via open-set text queries. In two prominent applications, zero-shot CBMs and decoding of visual features into natural language, we achieve state-of-the-art results. Our method can be applied to various other applications that were previously restricted to CLIP models. Our work thus removes this restriction, being applicable to any visual classifier. Our work also enables concept naming in Sparse Autoencoders (Bricken et al., 2023) trained on any visual classifier, using the technique from (Rao et al., 2024). Finally, as with any research work, this study has its own limitations which should be clearly acknowledged. We discuss limitations of our method in Appendix Section A.
Appendix
Appendix A Limitations and Ethical Concerns
As with any research work, our study has its own set of limitations that should be transparent and acknowledged. In particular, we identify two primary limitations of our method: the first concerns wrong semantic associations of class names in CBMs, while the second pertains to the limited generalization of our method to fine-grained datasets.
We start by addressing the first limitation of wrong semantic associations of class names in CBMs. This issue is closely tied to the choice of concept set used. When using the 20K most common words in English as our concept set, we observe that class names get associated to either wrong semantically-related concepts, or with components of the class name itself, making the detected concepts in CBMs less meaningful. Figure 8 illustrates such cases. In the first example, the top-detected concepts for the prediction “african grey” are concepts that are direct components of the class name itself (grey, african), variations of these words (gray, africa), or incorrect semantic associations with the word “african” that do not pertain to the bird itself (Ethiopian, Tanzania). Similarly, in the second example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric). This occurs because the bird “drake” is less familiar to the text encoder than the artist “drake”. In fact, a google search with the word “drake” yields directly the artist rather than the bird. In the third example, a similar issue arises with the animal “cock”, leading to associations with male reproductive terms. For each example, we also report the total logit score of the prediction.
However, it is worth noting the following:
  1. 1.
    Meaningful concepts such as “duck” (second example) and “hen” (third example) are still detected among the top concepts.
  2. 2.
    The incorrect semantic associations contribute only a negligible portion of the total logit, accounting for approximately 0.01% of the overall prediction score.
  3. 3.
    This issue is considerably less severe when using alternative concept sets, such as the LF-CBM concept set tailored for ImageNet.
  4. 4.
    This issue also appears in CLIP-based CBMs and hence not unique to our approach.
Figure 8: Limitations of our method is wrong semantic concept association in CBMs
The second limitation is the lack of generalization to fine-grained datasets. We test our ImageNet-trained reformulated classifiers to perform zero-shot transfer to the Places365 dataset (Zhou et al., 2017) for scene classification. It is important to note that Places365 is known to induce low performance, even when using supervised training with the powerful CLIP model, making it a challenging classification dataset. To perform zero-shot transfer, we convert the class names of Places365 into text prompts of the form: “a scene of a {class} location” which are then encoded using the text encoder $T$ to obtain the classification weights $U_{places}$. The image is encoded with the respective vision encoder to obtain the visual features $f$, which are then projected via the MLP to yield $\tilde{f}$. Zero-shot classification is performed using $\text{argmax}(\tilde{f}.U^{T}_{places})$. Results on the Places365 validation set are shown in Table 5. We report the supervised learning performance by training a linear probe on top of the CLIP model features. The supervised top-1 accuracy is 53.40 and 55.10 for ResNet50 and ViT-B/16 CLIP models, respectively. As demonstrated, these low top-1 accuracies suggest that Places365 is a particularly challenging dataset for classification. We then report the zero-shot performance of our method on several visual classifiers. The best-performing model achieves a top-1 zero-shot accuracy of 14.36%, much lower than the supervised performance. However, we observe that Transformer models based on large-scale pretraining perform better, a trend that aligns with findings in the out-of-distribution (OOD) literature. Moreover, our approach performs significantly better than a randomly shuffled classifier where we shuffle its weights across classes.
Model Top-1 (%) Top-5 (%)
Supervised
CLIP RN50 Linear Probe 53.40 -
CLIP ViT-B Linear Probe 55.10 -
Zero-Shot
Shuffled Classifier 0.24 1.54
DenseNet161 10.55 25.18
ResNet50 10.58 25.14
ResNet101 10.97 25.42
ViT-B/16 11.47 26.53
ResNet50v2 11.66 27.40
ConvNeXt-Base 11.76 27.12
Swin-Base 12.12 26.83
Swinv2-Base 12.13 27.18
EfficientNetv2-M 12.31 28.33
ViT-B/16v2 13.11 29.54
DINOv2-Base 13.96 32.35
ConvNeXtv2 (pt) 14.05 30.92
BeiT-L/16 14.15 31.08
ViT-B/16 (pt) 14.22 32.18
ConvNeXt (pt) 14.36 31.64
Table 5: Top-1 and Top-5 zero-shot transfer results to Places365 for various models
Appendix B Ablation Studies on the Text Encoder
In this section, we present ablation studies using other text encoders. We test a variety of text encoders from the Sentence-BERT library (Reimers & Gurevych, 2019) using the ResNet50 visual classifier. Results are presented in Table 6. We observe that the choice of the text encoder has minimal effect on the performance, as even lower-performing text encoders are capable of understanding class names.
Text Encoder Top-1 (%)
DistilRoberta 75.73
MPNet-Base 75.78
MPNet-Base-MultiQA 75.76
MiniLM 75.80
Table 6: Ablation studies on other text encoders
Appendix C Performance on Additional Models
We report performance on additional models that were not included in the main manuscript in Table 7.
Model Top-1 Orig. $\Delta$
ConvNeXt-Tiny 82.19 82.52 --0.33
ViT-B/32 75.40 75.91 --0.51
Swin-Small 82.63 83.20 --0.57
Swinv2-Tiny 81.44 82.07 --0.63
CvT-21 80.45 81.27 --0.82
[…]
Swinv2-Small 83.32 83.71 --0.39
[…]
Table 7: Performance of our reformulated classifiers for additional models
Appendix D Implementation Details
For the text encoder, we use the all-MiniLM-L12-v1555https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (Reimers & Gurevych, 2019). This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs. It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba et al., 2016), a GELU activation function (Hendrycks & Gimpel, 2016) and Dropout (Srivastava et al., 2014) with a drop probability of 0.5. The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function. The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder). We train the MLP projector with a batch size of 256 using the ADAM optimizer (Kingma & Ba, 2015) with a learning rate of 1e-4 that decays using a cosine schedule (Loshchilov & Hutter, 2017) over the total number of epochs. We follow the original image sizes that the classifier was trained on.
For the training images, we apply the standard image transformations that all classifiers were trained on which include a Random Resized Crop and a Random Horizontal Flip. For the validation images, we follow exactly the transformations that the classifier was evaluated on, which include resizing the image followed by a Center Crop to the image size that the classifier expects. Each model is trained on a single NVIDIA GeForce RTX 2080 Ti GPU.
Appendix E Process of Decoding Visual Features
We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$. A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$. To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li & Liang, 2021), which prepends learnable tokens in the embedding space. We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly. Our method builds upon the work of (Tewel et al., 2021).
A high-level overview of this process is illustrated in Figure 9. Using a pre-trained language model $G$, we prepend randomly initialized learnable tokens, referred to as prefixes, which guide $G$ to produce text that maximizes alignment with visual features. These learnable prefixes function as key-value pairs in each attention block, ensuring that every generated word can attend to them.
For each iteration $j$, at a timestep $ts$, we sample the top-$Q$ tokens from the output distribution of $G$, denoted as $G_{out}$, which serve as possible continuations for the sentence. These $Q$ candidate sentences are then encoded by a text encoder $T$, mapping them into the same embedding space as $\tilde{f}$. We compute the cosine similarity between each encoded sentence and $\tilde{f}$, resulting in $Q$ similarity scores. These scores are normalized with softmax and define a target distribution used to train $G_{out}$ via Cross-Entropy loss. The learnable prefixes are updated through backpropagation.
With the updated prefixes, $G$ is run again, and the most probable token is selected as the next word. This process is repeated for a predefined number of timesteps (up to the desired sentence length) or until the <.><.>< . > token is generated. At the end of each iteration, a full sentence is generated. We conduct this process for 20 iterations, generating 20 sentences in total. The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
We also add the fluency loss from (Tewel et al., 2021) as well as other token processing operations. We refer readers to (Tewel et al., 2021) for more information. We use the small GPT-2 of 124M parameters as $G$. We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
Figure 9: The process used to decode visual features of an image. The process is shown for the first timestep $ts=1$ with a hard prompt set as “an image of a”.
Appendix F Using only class names
We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP. In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin, 1998) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt et al., 2023) or VCVD(Menon & Vondrick, 2023). However, this approach would be considered as “cheating,” as it fails to faithfully replicate the classifier. The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes. Explicitly providing additional information would not replicate the classifier faithfully, and would also force the classifier to focus on predefined features or those we intend it to learn. Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation. For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts. For these reasons, we refrain from using any other additional information than the class names. We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
Appendix G Compositional Captioning
The prompt we used for compositional captioning is as follows:
I will give you several attributes and verbs that are included in an image, each with a score. The score reflects how important (or how grounded) the attrribute/verb is to the image, and higher means more important and grounded. Your job is to formulate a caption that describes the images by looking at the attributes/verbs with their associated scores. You should not reason or generate anything that is based on your own knowledge or guess. Everything you say has to be grounded in the attrbiutes/verbs and score importances. Please use the following the structure, style, and pattern of the following examples. Example 1: A woman wearing a net on her head cutting a cake. Example 2: A child holding a flowered umbrella and petting a yak. Example 3: A young boy standing in front of a computer keyboard. Example 4: a boy wearing headphones using one computer in a long row of computers. Example 5: A kitchen with a stove, microwave and refrigerator. Example 6: A chef carrying a large pan inside of a kitchen. Here are the attributes and scores: {detected concepts with scores}, and these are the verbs and scores: {detected verbs with scores}.
Next, we explore alternative concept sets in Compositional Captioning. In the main manuscript, we reported results using the 20,000 most common English words as our concept set. Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative. This allows for the generation of captions (here, decoded visual features) tailored to a specific domain. Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset. Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset. We use the ImageNet-specific concept set from (Oikarinen et al., 2023) and report zero-shot captioning performance in Table 8. As shown, we observe a decrease in all metrics. This shows that our method can readily decode visual features into text for any domain. Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
Method B4 M R-L C S
ZeroCap 2.6 11.5 14.6 5.5
ConZIC 1.3 11.5 12.8 5.2
Ours
MobileNetv3-L 3.50 12.7 29.1 11.4 6.1
ResNet50 3.60 12.7 29.3 12.0 6.0
ResNet101v2 3.50 12.9 29.3 12.2 6.3
WideResNet101v2 3.70 12.9 29.6 12.4 6.2
ConvNeXt-Base 3.80 12.8 29.5 12.7 6.2
EfficientNetv2-S 3.70 12.9 29.6 12.9 6.3
ViT-B/16 (pt) 3.80 13.1 29.5 13.2 6.5
BeiT-L/16 3.90 13.2 29.6 13.4 6.6
Table 8: Composition Captioning Performance using the ImageNet-specific LF-CBM concept set

### v2

@@
-Unlocking Open-Set Language Accessibility in Vision Models
+Unlocking Text Capabilities in Vision Models
 Abstract
-Visual classifiers offer high-dimensional feature representations that are challenging to interpret and analyze.
+Visual classifiers provide high-dimensional feature representations that are challenging to interpret and analyze.
 Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior.
-We propose a simple, yet powerful method for reformulating any visual classifier so that it can be accessed with open-set text queries without compromising its original performance.
-Our approach is label-free, efficient, and preserves the underlying classifier’s distribution and reasoning processes.
-We thus unlock several text-based interpretability applications for any classifier.
-We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language.
-In both applications, we achieve state-of-the-art results, greatly outperforming existing works.
-Our method enables text approaches for interpreting visual classifiers.
-Machine Learning, ICML
+We propose a simple, yet powerful method for reformulating any pretrained visual classifier so that it can be queried with free-form text without compromising its original performance.
+Our approach is label-free, data and compute-efficient, and is trained to preserve the underlying classifier’s distribution and decision-making processes.
+Our method unlocks several zero-shot text interpretability applications for any visual classifier.
+We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any visual classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language sentences.
+In both tasks we establish new state-of-the-art results, outperforming existing works and surpassing CLIP-based baselines with ImageNet-only trained classifiers, while using up to 400$\times$ fewer images and 400,000$\times$ less text during training.
 1 Introduction
-Visual classifiers inherently and purely provide visual signals, which are high-dimensional dense vector representations of images that are hard to interpret and dissect, limiting attribution maps as the main medium for interpretation.
-Attribution maps provide general and high-level information, indicating where in the image the most important region is for the prediction.
-Nevertheless, text is significantly more expressive, human-friendly, specific, and unlocks a variety of other interpretability applications that require language (Tewel et al., 2021; Koh et al., 2020; Oikarinen & Weng, 2023; Song et al., 2022; Minderer et al., 2022).
-Therefore, a question that arises is how can we talk or access those visual classifiers via open-set text?
+Visual classifiers provide dense, high-dimensional visual feature vectors that are difficult to interpret directly.
+Attribution maps—the standard way to interpret visual classifiers—only highlight the image region most influential to a prediction, offering a high-level spatial insight.
+They reveal where the key features lie, but not what those features actually mean or represent.
+By contrast, text is significantly more expressive, more precise and human-friendly, and it enables a new suite of text interpretability tasks  [1, 2, 3].
+This raises a key question: How can we query a purely visual classifier with free-form text?
 Performing this allows us to have a text interface that can be used to query visual classifiers.
-So far, this has only been achieved with generic vision-language models such as CLIP-based ones (Radford et al., 2021), which learn a shared vision-language space and thus enables any text to query the visual model directly.
-However, the deep learning field extends far beyond CLIP.
-For example, industries—being the primary end-users of such models—often rely on custom visual models tailored to their specific needs.
-This creates an application bottleneck: any task involving both visual and textual interpretations becomes restricted to CLIP or similar approaches, limiting flexibility, forcing reliance on such models, and inheriting their limitations.
-Figure 1: We formulate any classifier to allow querying it and measuring its association with any text, while maintaining its performance and distribution. $f$ is the feature vector of an image.
-In this work, we lift this restriction by proposing a method to reformulate any visual classifier such that it can be accessed via open-set text (Figure 1).
+Contrastive Language-Image Pre-Training (CLIP)-based vision–language models answer this by learning a shared embedding space [4], enabling any text to query the visual model directly.
+Yet in many real-world scenarios a high-performing, task-specific legacy model already exists and typically outperforms zero-shot vision-language models [5].
+Retraining such a specialist on a large image–text corpus following the CLIP approach or the GPT-like approach is impractical, in terms of computational cost and need of a huge amount of image-text data.
+Moreover, this approach alters the decision-making process and distribution of the legacy model.
+As a result, most text-based interpretability techniques remain confined to CLIP-like models, leaving specialist vision models without a language-driven interface.
+Figure 1: (a) Text interpretability is typically limited to models trained to share a joint embedding space such as CLIP-based ones.
+(b) Our approach removes this constraint, enabling any visual classifier to be queried through any arbitrary (unaligned) text encoder in a highly data- and compute-efficient, label-free manner, while preserving the classifier’s original distribution and reasoning process and thus offering text-based interpretability of the original classifier.
+The values here represent the cosine similarity between the text and the visual features $f$.
+In this work, we lift this limitation by proposing a method to reformulate any frozen visual classifier such that it can be accessed via free-form text (Figure 1).
 Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier.
+The number of data points is also significantly reduced compared to CLIP-based approaches.
 Secondly, it is label-free, no labels are required to achieve this formulation.
-Thirdly, it is faithful to the classifier and preserves its underlying distribution and reasoning process.
+Thirdly, it is explicitly trained to preserve the original distribution and reasoning process of the classifier.
 Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
 A visual classifier assigns an image to a specific category from a predefined set of discrete class labels.
-In ImageNet-1K (Deng et al., 2009), there are a total of $1,000$ labels.
+In ImageNet-1K [6], this set contains $1,000$ class labels.
 Originally, these discrete class labels correspond to class names in text format.
@@
 However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information.
-Specifically, these embeddings also semantically encode words that are close to the class name within the text embedding space.
-For the “goldfish” example, the closest words could be “freshwater”, “fins” and “orange”.
-Our method learns to map images into this text embedding space, thus associating both the class name and its surrounding semantic space with the image.
-This is accomplished through a trainable MLP that projects the visual features into the text embedding space, and is trained by aligning its output distribution across all classes with the original classifier’s output distribution, while keeping both the visual and textual encoders frozen.
+Specifically, these embeddings reside in a continuous space where nearby vectors capture related concepts.
+For the “goldfish" example, neighboring vectors might include “freshwater", “fins" and “orange".
+Our method learns to map images into this text embedding space using only class names, thus associating both the class name and its surrounding semantic space with the image.
+This is accomplished through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to output the same distribution scores across all classes, as the original classifier’s class distribution.
+This is done while while keeping both the visual and textual encoders frozen.
 By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space.
@@
 We demonstrate the effectiveness of our method with two primary applications.
-1) We build label-free concept bottleneck models for any classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the new linear classifier remains faithful to the original classifier.
+1) We build label-free concept bottleneck models for any pretrained classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the newly built linear classifier shares the same reasoning process with the original classifier.
 2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
-In summary, our contributions are as follows:
-  • We propose an efficient, label-free method to reconfigure any visual classifier such that it can be queried with open-set text queries, while also being faithful to the classifier.
-  • We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free, faithful zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language.
-  • Our method sets new state-of-the-art results outperforming existing works, including those which use CLIP as the visual encoder, although being trained with 400$\times$ less images and 400,000$\times$ less captions, thus efficiently learning faithful text interfaces to interpret any visual classifier.
+In summary, our contributions are as follows: (i) We propose an efficient, label-free method to reformulate any visual classifier such that it can be queried with open-set text queries, maintaining its performance within a negligible margin (average of 0.2 points drop in accuracy) (ii) We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language.
+(iii) On both these applications, our method sets new state-of-the-art results outperforming existing works, including those using CLIP as the visual encoder, while being trained with 400$\times$ less images and 400,000$\times$ less text data.
 2 Related Work
-While CLIP is the predominant approach to interpret vision models through language (Oikarinen & Weng, 2023), there exist works that try to decode visual features of models beyond CLIP.
-DeVIL (Dani et al., 2023) trains an autoregressive text generator to map visual features from different layers of a classifier into image captions, leveraging annotated image-caption pairs as ground-truth data.
-Similarly, Natural Language Explanations (NLEs) (Park et al., 2018; Kayser et al., 2021; Sammani et al., 2022) use annotated textual explanations in place of conventional captions.
-Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, and are therefore not faithful to the classifier.
-ZS-A2T (Salewski et al., 2023) maps attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective), and thus cannot be applied to visual classifiers.
-Our work on the other hand decodes visual features of any classifier in a zero-shot manner without requiring any annotated data, while also maintaining faithfulness to the original classifier.
-Text-to-Concept (T2C) (Moayeri et al., 2023) is the closest related work to ours, where a linear layer is trained to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
-Since the linear layer maps features to the CLIP space, T2C is strongly biased towards interpreting the CLIP model rather than the original classifier.
+While CLIP is the predominant approach to interpret vision models through language [2], there exist works that try to decode visual features of models beyond CLIP.
+DeVIL [7] and LIMBER [8] trains an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data.
+Similarly, Natural Language Explanations (NLEs) [9, 10, 11] use annotated textual explanations in place of conventional captions.
+Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process.
+ZS-A2T [12] converts attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective); therefore, it cannot be applied to standalone visual classifiers.
+In contrast, our method decodes visual features of any pretrained classifier in a zero-shot manner without requiring any annotated data, and is explicitly trained to keep the model’s predictive distribution intact.
+Text-to-Concept (T2C) [13] and [14] are the closest related works to ours.
+These works train a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
+Since the linear layer maps features to the CLIP space, these works are strongly biased towards interpreting the CLIP model rather than the original classifier.
 As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image.
-Moreover, T2C relies on CLIP supervision.
-In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution, ensuring faithfulness to it.
+Moreover, these works rely on CLIP supervision during training.
+In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution.
 3 Method
@@
 (a) The process of training the MLP mapping between vision and text space.
-(b) The process of inference with adapted visual classifier.
+A PyTorch-like pseudocode for (a) in included in Section B of the Appendix.
+(b) The process of inference with the adapted visual classifier.
 The text encoder acts as a linear classifier weight generator in our re-formulated classification process.
@@
 Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
-We propose to learn a simple light-weight MLP mapping function which projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$.
+We propose to learn a simple lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$.
 That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$.
 Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient.
-We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}” where {class} represents the class name in text format.
+We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}" where {class} represents the class name in text format.
 This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$.
 Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$.
+Here, $U$ acts as weights of the classification layer for our approach.
 We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$.
-Equivalently, this can be performed as a single matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class.
-The most straightforward approach to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution.
-However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions).
+Equivalently, this can be performed as a single vector-matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class.
+In other words, $S$ represents the classification logits of our approach.
+The most straightforward way to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution.
+However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions) and sacrificing text interpretability of the original classifier.
 Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored.
@@
 This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification.
-This makes our approach label-free and preserves the distribution and reasoning process of the classifier $F$.
-We provide a PyTorch-like pseudocode of our training approach in Listing LABEL:lst:pseudo, and illustrate it in Figure 2(a).
-It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions or hierarchies (see Section F in the Appendix for more information).
+Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters.
+Eq.
+1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose.
+Note also that this loss function naturally encodes the classifier’s relationship across all classes.
+Our training approach is illustrated in Figure 2(a), and we provide a PyTorch-like pseudocode of it in Section B of the Appendix.
+It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section I in the Appendix for more information).
 After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features.
@@
 This is shown in Figure 2(b).
-We show later that this weight generating process is important to build faithful concept bottleneck models.
-# text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
-# classifier: linear classifier weights of a frozen vision_encoder, shape (visual_dim, num_classes)
-# mlp: trainable MLP from visual_dim -> text_dim
-# images: batch of B images, shape (N, 3, height, width)
-visual_feats = vision_encoder(images) # (N, visual_dim)
-logits = visual_feats @ classifier # (N, num_classes)
-original_dist = softmax(logits, dim=-1) # (N, num_classes)
-mapped_feats = mlp(visual_feats) # (N, text_dim)
-mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
-text_feats = l2_norm(text_feats) # (N, text_dim)
-pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
-pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
-# cross entropy with original model’s soft distribution
-loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
-loss.backward() # only mapper parameters are updated
-Listing 1: PyTorch-like pseudocode for our method
 4 Applications
@@
 4.1 Zero-Shot Concept Bottleneck Models
-Concept Bottleneck Models (CBMs) (Koh et al., 2020) are a class of inherently interpretable models and have recently attracted significant attention.
+Concept Bottleneck Models (CBMs) [1] are a class of inherently interpretable models and have recently attracted significant attention.
 They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions.
-In (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score which represents the concept activation to the image.
-In (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of dense features.
-Recently, the Label-Free CBMs (LF-CBMs) family uses CLIP to perform the concept discovery step without annotated image-concept data, by either using CLIP as a model to provide ground-truth image-concept similarities to train a layer which performs concept discovery (Oikarinen et al., 2023), or by querying the image features from a set of predefined concepts within the CLIP space (Yang et al., 2022; Panousis et al., 2023) and using the cosine similarity scores between the concepts and the image as concept activations.
-However, there are major issues in the current literature of CBMs.
+In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image.
+In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of opaque features.
+Recently, Label-Free CBMs (LF-CBMs) use CLIP to perform the concept discovery step without annotated image-concept data.
+This is achieved either by using CLIP as a model to provide ground-truth image-concept similarities to train a layer that performs concept discovery [15], or by querying the image features from a set of predefined concepts within the CLIP space [16, 17] and using the cosine similarity scores between the concepts and the image as concept activations.
+However, there are major limitations in this existing line of CBMs.
 Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model.
-Secondly, all CBM approaches to date require training a linear probe on the concept activations which 1) hinders an on-the-fly ready-to-deploy CBMs, and 2) disconnects $W^{con}$ from the original feature-based classifier $U$, since $W^{con}$ is trained from scratch in a different space using ground-truth data and is hence not faithful to the original model, and 3) requires training a new $W^{con}$ classifier for every different set of concepts, hindering flexibility to the choice of concepts.
-Our method allows us to solve all the aforementioned issues and formulate zero-shot CBMs for any classifier.
-We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ represents the classification weights of the newly formulated classifier (the output of the text encoder $T$ for the class prompts).
+Secondly, all CBM approaches to date require training a linear probe on the concept activations, which 1) hinders an on-the-fly ready-to-deploy CBM at test-time directly, and 2) requires training a new $W^{con}$ classifier from scratch for every different set of concepts, hindering flexibility to the choice of concepts.
+Our method solves all the aforementioned issues, formulating zero-shot CBMs for any pretrained classifier.
+We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier.
 We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$.
-Following other works (Rao et al., 2024), and without loss of generality, we use the $Z=20K$ most common words in English [^2] https://github.com/first20hours/google-10000-english as our concept set.
-These are general concepts that represent world knowledge and are not tailored towards any specific dataset.
+Following other works [18], and without loss of generality, we use the $Z=20K$ most common words in English [19] as our concept set.
+These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset.
+To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set.
+Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”).
+In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
+This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
 We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$.
@@
 By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$.
-For an image $I$, we extract its visual features $f$ and use MLP to map them to $\tilde{f}$ which now lies in the text embedding space.
+For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space.
 That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$.
 Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to.
-That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and $C$.
+That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$.
 The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts.
-We provide an illustration in Figure 3.
-Figure 3: Our proposed approach to concept discovery in CBMs.
+We provide an illustration in Figure 3(a).
+Figure 3: (a) Our approach to concept discovery in CBMs.
+(b) Decoding visual features into natural language using any pretrained language decoder (e.g., GPT-2).
+We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features.
+Shown for iteration $j=1$
 We now show how to build the classifier $W^{con}$ that takes in the concept activations and outputs a distribution $S_{cn}$ over classes.
@@
 […]
-By observing Eq.
-2, we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
+From Eq.
+2 we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
 Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$.
@@
 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts.
-This means our CBMs are faithful to the original classifier.
-By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner, for any classifier, while also maintaining faithfulness to the original classifier and being flexible in the concept set.
+This means our CBMs preserve the basis reasoning process of the original classifier.
+By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner and for any classifier.
+Note that our method is flexible in the concept set.
+The concept set is chosen at test time directly, and the CBM is built directly on-the-fly.
 4.2 Zero-Shot Decoding of Visual Features into Text
-In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into natural language sentence.
+In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into a natural language sentence.
 This offers a text-based interpretation of what the visual features contain.
-We adapt the method introduced in (Tewel et al., 2021) for our purpose.
+We adapt the method introduced in [3] for our purpose.
 Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$.
@@
 Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text.
-We utilize an off-the-self pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text.
-We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li & Liang, 2021) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
-Specifically, we attach a set of learnable “virtual” tokens to $G$.
+We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text.
+We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [20] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
+Specifically, we attach a set of learnable “virtual" tokens to $G$.
 Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number.
@@
 We perform this process for several iterations.
-An overview of this process is shown in Figure 4.
-As this process is not the core contribution of our work, we leave more details to Section E of the appendix, which also includes a detailed illustration in Figure 9.
-Figure 4: To decode visual features into natural language using any pretrained language decoder (e.g., GPT-2), we apply prefix tuning while keeping the language decoder fixed, generating text that maximizes the similarity with the visual features.
-The figure is shown for iteration $j=1$
+An overview of this process is shown in Figure 3(b).
+As this process is not the core contribution of our work, we leave more details to Section H of the appendix, which also includes a detailed illustration in Figure 2.
 5 Experiments
 We first provide results on our newly formulated visual classifiers.
-We use the ImageNet-1K dataset due to the widespread availability of visual classifiers trained and evaluated on it.
+We use the challenging ImageNet-1K dataset due to the widespread publicly available visual classifiers trained and evaluated on it.
 We apply our method on a diverse set of 40 visual classifiers.
-For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He et al., 2015), Wide ResNets (Zagoruyko & Komodakis, 2016), ResNeXts (Xie et al., 2016), ShuffleNetv2 (Ma et al., 2018), EfficientNetv2 (Tan & Le, 2021), Densely Connected Networks (DenseNets) (Huang et al., 2016), ConvNeXts (Liu et al., 2022) and ConvNeXtv2 (Woo et al., 2023).
-For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (Dosovitskiy et al., 2021), DINOv2 (Oquab et al., 2024), BeiT (Bao et al., 2022), the hybrid Convolution-Vision Transformer CvT (Wu et al., 2021), Swin Transformer (Liu et al., 2021b) and Swin Transformer v2 (Liu et al., 2021a).
-All models are pretrained on ImageNet-1K from the PyTorch [^3] https://pytorch.org/vision/stable/models.html and HuggingFace (Wolf et al., 2020) libraries.
+For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [21], Wide ResNets [22], ResNeXts [23], ShuffleNetv2 [24], EfficientNetv2 [25], Densely Connected Networks (DenseNets) [26], ConvNeXts [27] and ConvNeXtv2 [28].
+For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [29], DINOv2 [30], BeiT [31], the hybrid Convolution-Vision Transformer CvT [32], Swin Transformer [33] and Swin Transformer v2 [34].
+All models are pretrained on ImageNet-1K from the PyTorch [35] and HuggingFace [36] libraries.
 Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K.
-Models with a subscript v2 are trained following the updated PyTorch training recipe (Vryniotis, 2021).
+Models with a subscript v2 are trained following the updated PyTorch training recipe [37].
 Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k.
@@
 Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K.
-For the text encoder, we use the MiniLM Sentence Encoder (Wang et al., 2020), with ablation studies on alternative text encoders shown in Section B of the Appendix.
-For implementation details, we refer to Section D of the Appendix.
+For the text encoder, we use the MiniLM Sentence Encoder [38], with ablation studies on other text encoders presented in Section D of the Appendix.
 Results are presented in Table 1.
@@
 As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models.
-We refer to Appendix Section C for results on other models.
-Model Top-1 Orig. $\Delta$
-ResNet50 75.80 76.13 --0.33
+We also perform ablation studies on the MLP to verify its effectiveness in Section C of the Appendix.
+We refer to Appendix Section F for results on other models and to Section G for implementation details.
+Model Top-1 Orig. $\Delta$ Model Top-1 Orig. $\Delta$
+ResNet50 75.80 76.13 --0.33 ResNeXt50-32x4dv2 80.79 80.88 --0.09
 […]
 […]
-ResNet101 77.19 77.37 --0.18
-WideResnet50 78.35 78.47 --0.12
+ResNet101 77.19 77.37 --0.18 ViT-B/16 80.70 81.07 --0.37
 […]
 […]
-DenseNet161 77.04 77.14 --0.10
-DenseNet169 75.46 75.60 --0.14
-EfficientNetv2-S 84.04 84.23 --0.19
-EfficientNetv2-M 84.95 85.11 --0.16
 […]
-ConvNeXt-Small 83.42 83.62 --0.20
-ConvNeXt-Base 83.88 84.06 --0.18
 […]
-ResNeXt50-32x4d 77.44 77.62 --0.18
+DenseNet169 75.46 75.60 --0.14 Swin-Base 83.22 83.58 --0.36
+EfficientNetv2-S 84.04 84.23 --0.19 Swinv2-Base 83.72 84.11 --0.39
+EfficientNetv2-M 84.95 85.11 --0.16 BeiT-B/16 84.54 85.06 --0.52
 […]
-ResNeXt101-64x4d 83.13 83.25 --0.12
-ResNeXt101-32x8d 79.10 79.31 --0.21
-ViT-B/16 80.70 81.07 --0.37
-[…]
-ViT-L/32 76.72 76.97 --0.25
-ViT-L/16 79.56 79.66 --0.10
-Swin-Base 83.22 83.58 --0.36
-Swinv2-Base 83.72 84.11 --0.39
-BeiT-B/16 84.54 85.06 --0.52
-BeiT-L/16 87.22 87.34 --0.12
-DINOv2-B 84.40 84.22 +0.18
-ConvNeXtV2-B 84.56 84.73 --0.17
+ConvNeXt-Small 83.42 83.62 --0.20 DINOv2-B 84.40 84.22 +0.18
+ConvNeXt-Base 83.88 84.06 --0.18 ConvNeXtV2-B 84.56 84.73 --0.17
 […]
@@
 denotes the original Top-1 accuracy.
-$\Delta$ represents the difference between them $\Delta$ = Top-1 - Orig.
+$\Delta$ represents the difference between them ($\Delta$ = Top-1 -- Orig).
 Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2.
-As baselines, we compare with LF-CBMs (Oikarinen et al., 2023), LaBo (Yang et al., 2022), CDM (Panousis et al., 2023) and the recent state-of-the-art DN-CBMs (Rao et al., 2024).
-All these methods are supervised and use CLIP-based models.
+As baselines, we compare with 9 methods: LF-CBMs [15], LaBo [16], CDM [17], DCLIP [39], DN-CBMs [18], DCBM [40], Multi-Modal (MM) CBMs [41], CF-CBM [42] and SALF-CBM [43].
+All these methods are supervised and use CLIP-based models either for computing the concept activations or as a backbone.
 Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art.
 Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples.
-The best results are obtained by the ConvNeXtv2 and ViT-L/16 models, which achieve a top-1 accuracy of $86.3$.
+The best results are obtained by the ConvNeXtv2 model, which achieve a top-1 accuracy of $86.4$.
 All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
-Method Model Top-1
+Figure 4: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction.
+Effectivenss of Concept Interventions: As a supplementary experiment, we also report concept intervention results on our ZS-CBMs in Section E of the Appendix, showing the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
 Supervised CBMs
-LF-CBM CLIP ResNet50 67.5
-LF-CBM CLIP ViT-B/16 75.4
-LaBo CLIP ResNet50 68.9
-LaBo CLIP ViT-B/16 78.9
-CDM CLIP ResNet50 72.2
-CDM CLIP ViT-B/16 79.3
-DCLIP CLIP ViT-B/16 68.0
-DN-CBM CLIP ResNet50 72.9
-DN-CBM CLIP ViT-B/16 79.5
+Method Model Top-1 Method Model Top-1
+LF-CBM CLIP ResNet50 67.5 DN-CBM CLIP ViT-B/16 79.5
+LF-CBM CLIP ViT-B/16 75.4 SALF-CBM-S CLIP ViT-B/16 75.3
+LaBo CLIP ResNet50 68.9 SALF-CBM-D CLIP ViT-B/16 76.3
+LaBo CLIP ViT-B/16 78.9 DCBM-SAM2 CLIP ViT-L/14 77.9
+CDM CLIP ResNet50 72.2 DCBM-GDINO CLIP ViT-L/14 77.4
+CDM CLIP ViT-B/16 79.3 DCBM-RCNN CLIP ViT-L/14 77.8
+DCLIP CLIP ViT-B/32 63.0 MM-CBM CLIP ViT-B/16 77.9
+DCLIP CLIP ViT-B/16 68.0 MM-ProtoSim CLIP ViT-B/16 78.8
+DCLIP CLIP ViT-L/14 75.0 CF-CBM-H CLIP ViT-B/16 77.4
+DN-CBM CLIP ResNet50 72.9 CF-CBM-L CLIP ViT-B/16 78.4
 Zero-Shot CBMs (Ours)
-ZS-CBM ResNet50 73.9
-[…]
-ZS-CBM ResNet101 75.3
-[…]
-ZS-CBM WideResNet50 76.9
+Method Model Top-1 Method Model Top-1
+ZS-CBM ResNet50 73.9 ZS-CBM ViT-B/32 73.3
 […]
 […]
-ZS-CBM DenseNet121 69.9
-ZS-CBM DenseNet161 75.2
-ZS-CBM EfficientNetv2-S 83.0
-ZS-CBM EfficientNetv2-M 83.9
-ZS-CBM ConvNeXt-Small 81.9
-ZS-CBM ConvNeXt-Base 82.8
-ZS-CBM ViT-B/32 73.3
-ZS-CBM ViT-B/16 79.3
 […]
-ZS-CBM Swin-Base 82.2
-ZS-CBM Swinv2-Base 82.6
-[…]
-ZS-CBM BeiT-B/16 83.0
-ZS-CBM DINOv2-B 82.6
+ZS-CBM WideResNet50 76.9 ZS-CBM Swinv2-Base 82.6
 […]
 […]
-ZS-CBM BeiT-L/16 86.2
+ZS-CBM DenseNet121 69.9 ZS-CBM DINOv2-B 82.6
 […]
 […]
-Table 2: Performance of Zero-Shot CBMs (ZS-CBM) on ImageNet validation set for several classifiers.
-Figure 5: Qualitative examples of our zero-shot CBMs.
-We show the top-detected concepts, each with their corresponding importance score to the prediction.
-All examples use the 20K concept set, except for the first which uses the ImageNet LF-CBM concept set
-Figure 6: Global class-wise interpretability analysis with our Concept Bottleneck Model.
-We highlight in yellow the top concepts in ”hammerhead shark” that are not present in ”tiger shark”, and therefore distinctive to ”hammerhead shark”.
+ZS-CBM EfficientNetv2-M 83.9 ZS-CBM BeiT-L/16 86.2
+[…]
+[…]
+Table 2: Performance of Supervised and Zero-Shot CBMs on the ImageNet validation set.
 In Figure 4, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance.
 The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class.
-By observing the second example, we see that the image is predicted as a “goose” because it has duck-like features, it is in the size of a swan, and it has bird and pigeon-like features.
-Similarly, we can see that the third image is a “spotted salamander” due to its lizard-like structure, a head that looks like the frog’s head, a snake-like tail and spots like a leopard.
-In the last image, we can see that the prediction is “ostrich” due to its camel-like hump, mammal-like size, piegeon-like body, the horse-riding-like feature, and is a type of a bird.
-In Figure 6, we present a probability distribution of global class-wise concepts.
-These are concepts detected for all images of a specific class, along with their frequency.
-We consider two semantically similar classes but distinctively different: “hammerhead shark” and a “tiger shark”.
-We highlight in yellow the top concepts in “hammerhead shark” that are not present in “tiger shark”.
-These concepts are “harpoon” and “lobster hammer”, both which are distinctive to the head of the hammerhead shark and drive its prediction.
-Figure 7: Qualitative examples of decoded visual features from different visual classifiers, into natural text.
+We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder.
+All examples use the LF-CBM concept set [15].
+By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab.
+Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [44] regarding the “dumbbell” class.
+Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it.
+With our method, we can obtain a textual interpretation of the biases that the original classifier learns.
+We also provide qualitative examples of global class-wise concepts detected in Section K of the appendix.
 Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features.
-Since the ground-truth content of visual features is unknown, we rely on COCO captions (Lin et al., 2014) for evaluation.
-While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose.
+Since the ground-truth content of visual features is unknown—that is, we do not know exactly what the model uses for reasoning—we rely on COCO captions [45] for evaluation.
+While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose, and represent the best available alternative for evaluation.
 Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person).
@@
 Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark.
-We present results on the widely used “Karpathy test split” with various vision classifiers.
-As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap (Tewel et al., 2021) and ConZIC (Zeng et al., 2023), both which use CLIP.
-For evaluation, we employ standard natural language generation (NLG) metrics: BLEU-4 (B@4) (Papineni et al., 2002), METEOR (M) (Banerjee & Lavie, 2005), ROUGE-L (R-L) (Lin, 2004), CIDEr (C) (Vedantam et al., 2014), and SPICE (S) (Anderson et al., 2016).
-Results are shown Table 3.
+We present results on the widely used “Karpathy test split" with various vision classifiers.
+As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap [3] and ConZIC [46], both which use CLIP.
+For evaluation, we employ standard natural language generation metrics: BLEU-4 (B@4) [47], METEOR (M) [48], ROUGE-L (R-L) [49], CIDEr (C) [50], and SPICE (S) [51].
+Results are shown in Table 3(a).
 ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems.
@@
 We also present qualitative examples of the decoded visual features for different models in Figure 5.
-We can see from the first example, that BeiT-L/16 contains both the vegetables and dog in its features, while ConvNexTv2 only focuses on the vegetables, and ViT-B/16 and DINOv2 only focus on on the dog.
+From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables.
+In contrast, ViT-B/16 and DINOv2 focus exclusively on the dog and its characteristics.
 Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason.
-This application allows a user-friendly way of interpreting visual features, applicable even to a layman user.
+This application allows human-friendly interpretations of visual features, understandable even to a layman user.
 Model B4 M R-L C S
@@
 ConZIC 1.3 11.5 12.8 5.2
+Ours
 DenseNet161 1.50 10.2 20.4 15.8 6.3
@@
 […]
-Table 3: Zero-Shot Image Captioning Performance
+(a) Zero-Shot Image Captioning Performance
 Model B4 M R-L C S
@@
 […]
+[…]
 ResNet101v2 4.30 12.6 30.1 18.0 6.8
-ConvNeXt-Base 4.50 12.7 30.3 18.1 6.9
+[…]
 […]
@@
 DINOv2-Base 4.60 13.0 30.7 18.7 7.1
-Table 4: Composition Captioning Performance across models.
-Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics.
+(b) Zero-Shot Compositional Captioning Performance
+Table 3: Comparison of Zero-Shot Captioning Performance
+Figure 5: Qualitative examples of decoded visual features from different visual classifiers, into text.
+Note that our results in Table 3(a) are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics.
 However, it is important to note that B4 and M are n-gram overlap-based metrics.
 They assume that the generated caption follows a specific structure and style.
-We verify this by revisiting an old image captioning paradigm termed as compositional captioning (Kulkarni et al., 2011), and later revived with deep learning methods (Lu et al., 2018).
+We verify this by revisiting an old image captioning paradigm termed as compositional captioning [52], and later revived with deep learning methods [53].
 In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence.
 With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner.
-Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3, and feed them, along with their similarity scores, to an LLM.
+Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3(a), and feed them, along with their similarity scores, to an LLM.
 For the concepts, we use the same concept set as in Section 4.1.
-We also add a list of the most common verbs [^4] https://github.com/datmt/English-Words-Updated in English to the pool.
+We also add a list of the most common verbs [54] in English to the pool.
 This allows us to cover all possible words and interactions.
@@
 This allows us to generate sentences adhering to a specific style and structure.
-For this experiment, we used GPT4o-mini (OpenAI, 2024) as our LLM, as it is fast and cost-efficient.
+For this experiment, we used GPT4o-mini [55] as our LLM, as it is fast and cost-efficient.
 In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances.
-We provide details on the exact prompt we used in Appendix Section G.
-Results are shown in Table 4.
-While results on CiDEr and SPICE are incremental compared to results in Table 3, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3 compared to baseline methods.
-Also note that since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating sentences in a specific style (e.g., humorous, positive, or negative).
-6 Conclusion and Future Directions
-In this work, we presented a method to transform visual classifiers such that they can be queried via open-set text queries.
-In two prominent applications, zero-shot CBMs and decoding of visual features into natural language, we achieve state-of-the-art results.
+We provide details on the exact prompt we used in Appendix Section J.
+Results are shown in Table 3(b).
+While results on CiDEr and SPICE are incremental compared to results in Table 3(a), the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3(a) compared to baseline methods.
+Also note that, since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains (see Section J.1 in the appendix).
+6 Conclusion
+We presented a method to transform visual classifiers such that they can be queried via open-set text queries.
+We achieve state-of-the-art results in two applications, i.e., zero-shot CBMs and decoding of visual features into natural language sentences.
 Our method can be applied to various other applications that were previously restricted to CLIP models.
-Our work thus removes this restriction, being applicable to any visual classifier.
-Our work also enables concept naming in Sparse Autoencoders (Bricken et al., 2023) trained on any visual classifier, using the technique from (Rao et al., 2024).
-Finally, as with any research work, this study has its own limitations which should be clearly acknowledged.
-We discuss limitations of our method in Appendix Section A.
-Appendix
-Appendix A Limitations and Ethical Concerns
-As with any research work, our study has its own set of limitations that should be transparent and acknowledged.
+Our work removes this restriction, being applicable to any visual classifier.
+Finally, as with any research work, this study has its own limitations which are discussed in Appendix Section A.
+Appendix A Limitations
+As with any research work, our study has its own limitations that should be transparent and acknowledged.
 In particular, we identify two primary limitations of our method: the first concerns wrong semantic associations of class names in CBMs, while the second pertains to the limited generalization of our method to fine-grained datasets.
@@
 This issue is closely tied to the choice of concept set used.
-When using the 20K most common words in English as our concept set, we observe that class names get associated to either wrong semantically-related concepts, or with components of the class name itself, making the detected concepts in CBMs less meaningful.
+When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts.
 Figure 1 illustrates such cases.
-In the first example, the top-detected concepts for the prediction “african grey” are concepts that are direct components of the class name itself (grey, african), variations of these words (gray, africa), or incorrect semantic associations with the word “african” that do not pertain to the bird itself (Ethiopian, Tanzania).
-Similarly, in the second example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric).
-This occurs because the bird “drake” is less familiar to the text encoder than the artist “drake”.
+In the first example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric).
+This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake".
 In fact, a google search with the word “drake” yields directly the artist rather than the bird.
-In the third example, a similar issue arises with the animal “cock”, leading to associations with male reproductive terms.
+In the second example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself.
+We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms).
 For each example, we also report the total logit score of the prediction.
@@
   1. 1.
-    Meaningful concepts such as “duck” (second example) and “hen” (third example) are still detected among the top concepts.
+    Meaningful concepts such as “duck" (first example) are still detected among the top concepts.
   2. 2.
@@
 The second limitation is the lack of generalization to fine-grained datasets.
-We test our ImageNet-trained reformulated classifiers to perform zero-shot transfer to the Places365 dataset (Zhou et al., 2017) for scene classification.
+We test our ImageNet-trained reformulated classifiers to perform zero-shot transfer to the Places365 dataset [56] for scene classification.
 It is important to note that Places365 is known to induce low performance, even when using supervised training with the powerful CLIP model, making it a challenging classification dataset.
-To perform zero-shot transfer, we convert the class names of Places365 into text prompts of the form: “a scene of a {class} location” which are then encoded using the text encoder $T$ to obtain the classification weights $U_{places}$.
+To perform zero-shot transfer, we convert the class names of Places365 into text prompts of the form: “a scene of a {class} location" which are then encoded using the text encoder $T$ to obtain the classification weights $U_{places}$.
 The image is encoded with the respective vision encoder to obtain the visual features $f$, which are then projected via the MLP to yield $\tilde{f}$.
@@
 The best-performing model achieves a top-1 zero-shot accuracy of 14.36%, much lower than the supervised performance.
-However, we observe that Transformer models based on large-scale pretraining perform better, a trend that aligns with findings in the out-of-distribution (OOD) literature.
-Moreover, our approach performs significantly better than a randomly shuffled classifier where we shuffle its weights across classes.
+However, we emphasize that the observed low generalization performance is not attributed to our formulation; rather, it stems from the inherent limitations of the classifier itself.
+If the classifier (e.g., ResNet101) fails to generalize to out-of-distribution (OOD) fine-grained classes—as is the case—then it is unreasonable to expect the text-transformed classifier to perform any better.
+Since our transformed classifier is a replication of the original classifier, it inherits all of its limitations.
+In fact, expecting improved performance from the transformed classifier would be counterintuitive, as it would indicate a departure from the original classifier’s underlying reasoning.
+Moreover, as demonstrated in Table 1, models that benefit from large-scale pretraining tend to achieve superior OOD fine-grained performance, a trend that is consistent with findings in the OOD literature [57].
+Finally, our approach performs significantly better than a randomly shuffled classifier where we shuffle its weights across classes.
 Model Top-1 (%) Top-5 (%)
@@
 Table 1: Top-1 and Top-5 zero-shot transfer results to Places365 for various models
-Appendix B Ablation Studies on the Text Encoder
-In this section, we present ablation studies using other text encoders.
-We test a variety of text encoders from the Sentence-BERT library (Reimers & Gurevych, 2019) using the ResNet50 visual classifier.
-Results are presented in Table 6.
-We observe that the choice of the text encoder has minimal effect on the performance, as even lower-performing text encoders are capable of understanding class names.
+Appendix B Pseudocode of our method
+We provide a PyTorch-like pseudocode of our training approach in Listing 1.
+# text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
+# classifier: linear classifier weights of a frozen vision_encoder, shape (visual_dim, num_classes)
+# mlp: trainable MLP from visual_dim -> text_dim
+# images: batch of B images, shape (N, 3, height, width)
+visual_feats = vision_encoder(images) # (N, visual_dim)
+logits = visual_feats @ classifier # (N, num_classes)
+original_dist = softmax(logits, dim=-1) # (N, num_classes)
+mapped_feats = mlp(visual_feats) # (N, text_dim)
+mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
+text_feats = l2_norm(text_feats) # (N, text_dim)
+pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
+pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
+# cross entropy with original model’s soft distribution
+loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
+loss.backward() # only mapper parameters are updated
+Listing 1: PyTorch-like pseudocode for our method
+Appendix C Ablation Studies of the MLP
+In this section, we evaluate the impact of the MLP and verify its effectiveness.
+We perform the following ablation studies: 1) Mean Ablation: For an image, we replace the input features to the MLP with a constant mean of the features calculated across the full ImageNet validation set.
+2) Random Features: For an image, we replace the input features to the MLP with random values sampled from a normal distribution with a mean and standard deviation equal to that of the features calculated across the full ImageNet validation set.
+3) Random Weight Ablation: We randomize the weights of the MLP projection.
+4) Shuffled Ablation: For an image, we replace the input features to the MLP with input features of another random image in the validation dataset.
+For each ablation experiment, we compute the ImageNet validation accuracy.
+We expect the accuracy to drop across all ablations.
+As shown in Table 2, the accuracy nearly drops to zero in every case, confirming the effectiveness of our MLP.
+Model Mean Feature $\downarrow$ Random Features $\downarrow$ Shuffled Features $\downarrow$ Random Weights $\downarrow$
+ResNet101v2 Ours 81.49 81.49 81.49 81.49
+Ablated 0.10 0.11 1.70 0.11
+ConvNeXt-Base Ours 83.88 83.88 83.88 83.88
+Ablated 0.10 0.11 1.79 0.10
+BeiT-L/16 Ours 87.22 87.22 87.22 87.22
+Ablated 0.10 0.11 1.87 0.11
+DINOv2-B Ours 84.40 84.40 84.40 84.40
+Ablated 0.10 0.13 1.76 0.09
+Table 2: Ablation studies of the MLP.
+Appendix D Ablation Studies on the Text Encoder
+In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library [58] with the ResNet50 visual classifier.
+We observe that the choice of the text encoder has minimal effect on the performance.
+This is because even lower-performing text encoders are capable of understanding class names.
 Text Encoder Top-1 (%)
@@
 Table 3: Ablation studies on other text encoders
-Appendix C Performance on Additional Models
+Appendix E Effectiveness of Concept Interventions
+As a supplementary experiment, we also report concept intervention results on our ZS-CBMs.
+Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
+The Waterbirds-100 dataset [59] is a standard dataset used in previous works [18] to conduct CBM intervention experiments.
+It is a binary classification dataset of two classes: waterbirds and landbirds.
+The training images of waterbirds are on water backgrounds, and training images of landbirds are on land backgrounds.
+However, the validation images do not have that correlation, where waterbird images appear on land backgrounds and landbird images on water backgrounds.
+The model is assumed to learn the water–land background correlation to perform this classification task.
+By building a CBM, we can correct this bias by intervening in concepts in the CB layer.
+However, there are two challenges associated with using the Waterbirds dataset in our work, and therefore we curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set.
+We provide more details about this process in Section L.
+Our curated dataset includes 140 validation images (70 for each class).
+We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from [18] which includes a collection of bird-related concepts and a collection of land-related concepts.
+The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification.
+To correct this, we intervene in the concepts in the CB layer, following the setup from [18].
+Intervention R (Int.
+R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop.
+Intervention K (Int.
+K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase.
+Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
+We also conduct concept intervention experiments with a more challenging multi-class setup.
+Unlike the Waterbirds dataset where bias-correlation issues are assumed, we make no such assumption here.
+We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy.
+A drop in accuracy indicates that these concepts are important for prediction.
+We select a subset of 10 classes from ImageNet following [60].
+We use the original ImageNet validation images for those classes, rather than the validation set from [60], because the latter contains original ImageNet training examples that our model has already seen.
+We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts.
+The classes and concepts we used are provided in Section M of the supplementary material.
+We then measure CBM accuracy before and after intervention.
+Results are provided in Table 5.
+For each image, we zero out the activations of its class-related concepts and report the Intervention R (Int.
+R) metric.
+As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
+Model Orig. CBM Int. (R)$\downarrow$ Int. (K)$\uparrow$
+BeiT-B/16 54.29 41.43 (-12.86) 58.57 (+4.28)
+[…]
+[…]
+DiNOv2 52.86 43.57 (-9.29) 59.29 (+6.43)
+BeiT-L/16 52.86 44.29 (-8.57) 58.57 (+5.71)
+Table 4: CBM Interventions on the standard Waterbirds dataset
+Model Orig. CBM Int. (R)$\downarrow$
+ResNet50 96.80 76.00 (-20.8)
+ResNet101 97.00 76.80 (-20.2)
+DINOv2-B 98.60 78.40 (-20.2)
+BeiT-B/16 98.60 78.80 (-19.8)
+[…]
+Table 5: CBM Interventions in a multi-class setup
+Appendix F Performance on Additional Models
 We report performance on additional models that were not included in the main manuscript in Table 6.
@@
 CvT-21 80.45 81.27 --0.82
-[…]
 Swinv2-Small 83.32 83.71 --0.39
@@
 Table 6: Performance of our reformulated classifiers for additional models
-Appendix D Implementation Details
-For the text encoder, we use the all-MiniLM-L12-v1555https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (Reimers & Gurevych, 2019).
+Appendix G Implementation Details
+For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library [58].
 This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs.
 It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
-Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba et al., 2016), a GELU activation function (Hendrycks & Gimpel, 2016) and Dropout (Srivastava et al., 2014) with a drop probability of 0.5.
+Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization [61], a GELU activation function [62] and Dropout [63] with a drop probability of 0.5.
 The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function.
 The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder).
-We train the MLP projector with a batch size of 256 using the ADAM optimizer (Kingma & Ba, 2015) with a learning rate of 1e-4 that decays using a cosine schedule (Loshchilov & Hutter, 2017) over the total number of epochs.
+We train the MLP projector with a batch size of 256 using the ADAM optimizer [64] with a learning rate of 1e-4 that decays using a cosine schedule [65] over the total number of epochs.
 We follow the original image sizes that the classifier was trained on.
@@
 Each model is trained on a single NVIDIA GeForce RTX 2080 Ti GPU.
-Appendix E Process of Decoding Visual Features
+Appendix H Process of Decoding Visual Features
 We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$.
 A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$.
-To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li & Liang, 2021), which prepends learnable tokens in the embedding space.
+To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning [20], which prepends learnable tokens in the embedding space.
 We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly.
-Our method builds upon the work of (Tewel et al., 2021).
+Our method builds upon the work of [3].
 A high-level overview of this process is illustrated in Figure 2.
@@
 The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
-We also add the fluency loss from (Tewel et al., 2021) as well as other token processing operations.
-We refer readers to (Tewel et al., 2021) for more information.
+We also add the fluency loss from [3] as well as other token processing operations.
+We refer readers to [3] for more information.
 We use the small GPT-2 of 124M parameters as $G$.
 We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
-Figure 9: The process used to decode visual features of an image. The process is shown for the first timestep $ts=1$ with a hard prompt set as “an image of a”.
-Appendix F Using only class names
+Figure 2: The process used to decode visual features of an image. The process is shown for the first timestep $ts=1$ with a hard prompt set as “an image of a".
+Appendix I Using only class names
 We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP.
-In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin, 1998) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt et al., 2023) or VCVD(Menon & Vondrick, 2023).
-However, this approach would be considered as “cheating,” as it fails to faithfully replicate the classifier.
+In practice, we can go beyond class names by using resources like a class hierarchy from WordNet [66] (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL [67] or VCVD[39].
+However, this approach would be considered as “cheating.
 The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes.
-Explicitly providing additional information would not replicate the classifier faithfully, and would also force the classifier to focus on predefined features or those we intend it to learn.
+Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn.
 Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation.
@@
 We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
-Appendix G Compositional Captioning
+Appendix J Compositional Captioning
 The prompt we used for compositional captioning is as follows:
@@
 Here are the attributes and scores: {detected concepts with scores}, and these are the verbs and scores: {detected verbs with scores}.
+J.1 Domain-Specific Compositional Captioning
 Next, we explore alternative concept sets in Compositional Captioning.
@@
 Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset.
-We use the ImageNet-specific concept set from (Oikarinen et al., 2023) and report zero-shot captioning performance in Table 8.
+We use the ImageNet-specific concept set from [15] and report zero-shot captioning performance in Table 7.
 As shown, we observe a decrease in all metrics.
@@
 Table 7: Composition Captioning Performance using the ImageNet-specific LF-CBM concept set
+Appendix K Global Class-Wise Qualitative Examples
+In Figure 3, we present a probability distribution of global class-wise concepts.
+These are concepts detected for all images of a specific class, along with their frequency.
+We consider two semantically similar classes but distinctively different: “hammerhead shark" and a “tiger shark".
+We highlight in yellow the top concepts in “hammerhead shark" that are not present in “tiger shark".
+These concepts are “harpoon" and “lobster hammer", both which are distinctive to the head of the hammerhead shark and drive its prediction.
+Figure 3: Global class-wise interpretability analysis with our Concept Bottleneck Model.
+We highlight in yellow the top concepts in "hammerhead shark" that are not present in "tiger shark", and therefore distinctive to "hammerhead shark".
+Appendix L Waterbirds/Landbirds Dataset
+There are two challenges associated with using the Waterbirds-100 dataset [59] in our work.
+First, our method cannot transform a classifier trained on this dataset because learning meaningful text representations from merely two labels (waterbird and landbird) with our MLP is not feasible.
+Second, we cannot evaluate our ImageNet-trained models on this dataset due to its highly artificial nature, which results in numerous severe out-of-distribution samples for classifiers trained on ImageNet.
+Therefore, we manually curated our own waterbird/landbird dataset using ImageNet validation images.
+Specifically, for the waterbird images, we consider classes of birds from ImageNet that are at least 90% of the time found in water backgrounds— a statistic verified by manually inspecting 100 random training images of those birds.
+For those birds, we then select their images in the ImageNet validation set that appear on land backgrounds.
+We perform a similar procedure for the landbird class.
+For the landbird class, we encountered a problem.
+Although many ImageNet bird classes nearly always feature land backgrounds in their training images, we were unable to find a substantial number of validation images depicting these birds against water backgrounds (e.g., we could not find any image of the bird robin with water background in the ImageNet validation set).
+In order to solve this issue, we utilized the Stable Diffusion 2.1 [68] text-to-image generative model to generate images of those birds on water backgrounds.
+We ensured that the generated images of those birds have the correct physical and distinctive features of the bird.
+For all images, we always ensure that the background (water/land) is clearly visible.
+This leads us to a validation dataset of 140 images (70 images for each class).
+Specifically, for the waterbird class, all 70 images coming from the ImageNet validation set.
+For the landbirds class, 15 images come from the ImageNet validation set, and 55 are generated with Stable Diffusion.
+Appendix M Multi-class CBM Intervention Classes and Concepts
+Class Concepts
+tench fish, freshwater, fins, dorsal, olive
+english springer dog, long ears, brown and white, playful, hunting
+cassette player portable, audio, tape, speakers, buttons
+chainsaw sharp, handheld, cutting, metal, wood
+church cross, tower, architecture, sacred, religious
+french horn curved, mouthpiece, musical instrument, orchestral, blow
+garbage truck large vehicle, wheels, clean, high load, lift
+gas pump fueling, hose, metallic, gasoline, handle
+golf ball small, white, round, rubber, dimples
+parachute fabric, fly, air, landing, strings
+Table 8: ImageNet classes and their five associated concepts we use in our multi-class CBM intervention experiment.

### v3

@@
-Unlocking Text Capabilities in Vision Models
+CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models
 Abstract
-Visual classifiers provide high-dimensional feature representations that are challenging to interpret and analyze.
-Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior.
-We propose a simple, yet powerful method for reformulating any pretrained visual classifier so that it can be queried with free-form text without compromising its original performance.
-Our approach is label-free, data and compute-efficient, and is trained to preserve the underlying classifier’s distribution and decision-making processes.
-Our method unlocks several zero-shot text interpretability applications for any visual classifier.
-We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any visual classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language sentences.
-In both tasks we establish new state-of-the-art results, outperforming existing works and surpassing CLIP-based baselines with ImageNet-only trained classifiers, while using up to 400$\times$ fewer images and 400,000$\times$ less text during training.
+Concept Bottleneck Models (CBMs) map dense, high-dimensional feature representations into a set of human-interpretable concepts which are then combined linearly to make a prediction.
+However, modern CBMs rely on the CLIP model to establish a mapping from dense feature representations to textual concepts, and it remains unclear how to design CBMs for models other than CLIP.
+Methods that do not use CLIP instead require manual, labor intensive annotation to associate feature representations with concepts.
+Furthermore, all CBMs necessitate training a linear classifier to map the extracted concepts to class labels.
+In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in a zero-shot manner.
+Our method is formulated by aligning the original classifier’s distribution (over discrete class indices) with its corresponding vision-language counterpart distribution derived from textual class names, while preserving the classifier’s performance.
+The approach requires no ground-truth image–class annotations, and is highly data-efficient and preserves the classifiers reasoning process.
+Applied and tested on over 40 visual classifiers, our resulting CLIP-free, zero-shot CBM sets a new state of the art, surpassing even supervised CLIP-based CBMs.
+Finally, we also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art results.
 1 Introduction
-Visual classifiers provide dense, high-dimensional visual feature vectors that are difficult to interpret directly.
-Attribution maps—the standard way to interpret visual classifiers—only highlight the image region most influential to a prediction, offering a high-level spatial insight.
-They reveal where the key features lie, but not what those features actually mean or represent.
-By contrast, text is significantly more expressive, more precise and human-friendly, and it enables a new suite of text interpretability tasks  [1, 2, 3].
-This raises a key question: How can we query a purely visual classifier with free-form text?
-Performing this allows us to have a text interface that can be used to query visual classifiers.
-Contrastive Language-Image Pre-Training (CLIP)-based vision–language models answer this by learning a shared embedding space [4], enabling any text to query the visual model directly.
-Yet in many real-world scenarios a high-performing, task-specific legacy model already exists and typically outperforms zero-shot vision-language models [5].
-Retraining such a specialist on a large image–text corpus following the CLIP approach or the GPT-like approach is impractical, in terms of computational cost and need of a huge amount of image-text data.
-Moreover, this approach alters the decision-making process and distribution of the legacy model.
-As a result, most text-based interpretability techniques remain confined to CLIP-like models, leaving specialist vision models without a language-driven interface.
-Figure 1: (a) Text interpretability is typically limited to models trained to share a joint embedding space such as CLIP-based ones.
-(b) Our approach removes this constraint, enabling any visual classifier to be queried through any arbitrary (unaligned) text encoder in a highly data- and compute-efficient, label-free manner, while preserving the classifier’s original distribution and reasoning process and thus offering text-based interpretability of the original classifier.
-The values here represent the cosine similarity between the text and the visual features $f$.
-In this work, we lift this limitation by proposing a method to reformulate any frozen visual classifier such that it can be accessed via free-form text (Figure 1).
-Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier.
+Visual classifiers predict a class as a linear combination of dense, high‑dimensional visual feature vectors that are difficult to interpret by humans.
+Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation).
+Predictions are then made as a linear combination of these concept activations.
+Initial CBMs required image–concept annotations to train the bottleneck layer that maps dense features to concepts.
+Modern CBMs (oikarinen2023labelfree; Yang2022LanguageIA; Panousis2023SparseLC; Rao2024DiscoverthenNameTC) overcome this limitation by leveraging the CLIP model (Radford2021LearningTV) as an annotation-free alternative.
+Since CLIP models map image and text into a shared embedding space, these approaches can query image features against a set of predefined textual concepts within that space and use cosine similarity scores to find a matching annotation.
+These approaches are commonly referred to as label-free CBMs.
+Nevertheless, in many real-world scenarios a high‑performing, task‑specific legacy model already exists and typically outperforms zero‑shot CLIP models (zhang2024why).
+A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models?
+Retraining such a specialist on a large image–text corpus following the CLIP approach is impractical, in terms of computational cost and need of a huge amount of image-text data.
+Furthermore, obtaining ground-truth image-concept annotations is time consuming and labor intensive.
+Finally, retraining this legacy model further alters its original decision-making process and distribution, which is typically not desired.
+As a result, CBMs remain confined to CLIP‑like models, leaving specialist vision models aside with no clear way to adapt them into the CBM framework.
+In this work, we lift this limitation by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP.
+TextUnlock has four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier.
 The number of data points is also significantly reduced compared to CLIP-based approaches.
 Secondly, it is label-free, no labels are required to achieve this formulation.
-Thirdly, it is explicitly trained to preserve the original distribution and reasoning process of the classifier.
+Thirdly, TextUnlock is trained to preserve the original distribution and reasoning process of the classifier, and does not compromise the classifier’s original performance (average of 0.2 points drop in accuracy).
 Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
+After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training but rather operating in a zero-shot manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
+In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP bottleneck (CLIP-free), neither annotated image-concept data (label-free) while also deriving the concept-to-class classifier in a zero-shot manner.
+(ii) We demonstrate the effectiveness of our method with 40 different architectures, along with extensive ablation studies and intervention results, and further show how our method can also be used to perform zero-shot image captioning.
+(iii) Our method sets new state-of-the-art results, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
+2 Related Work
+Concept Bottleneck Models (CBMs): Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) are a class of inherently interpretable models that map dense representations into human-understandable concepts, and then make predictions based on a linear combination of these interpretable units rather than opaque features.
+Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to perform the feature-to-concept mapping without requiring annotated image–concept data.
+This can be achieved either by using CLIP to provide ground-truth image–concept similarity scores that train the bottleneck layer (oikarinen2023labelfree), or by querying image features against a predefined set of concepts in the CLIP space (Yang2022LanguageIA; Panousis2023SparseLC) and using the resulting cosine similarities as concept activations.
+Several follow-up works (Rao2024DiscoverthenNameTC; Knab2024DCBMDV; Benou2025ShowAT; yuksekgonul2023posthoc) adopt this CLIP-based pipeline.
+However, these approaches remain limited to CLIP-based classifiers and cannot be applied to other visual models.
+Moreover, they require training a linear probe to map concept activations to class predictions.
+Our work addresses both limitations: we propose a fully CLIP-free CBM and demonstrate that our method can be readily used to derive the linear probe in a fully zero-shot manner.
+Transforming Visual Features to Text: There exist works that try to decode visual features of models beyond CLIP.
+DeVIL (Dani2023DeViLDV) and LIMBER (merullo2023linearly) train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data.
+Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process.
+ZS-A2T (Salewski2023ZeroshotTO) converts attention maps into natural language in a zero-shot manner using Large Language Models (LLMs), but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective).
+Text-to-Concept (T2C) (Moayeri2023TextToConceptB) also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
+Notably, all these methods rely on the CLIP approach and/or its supervision, and they alter the classifier’s distribution by entirely discarding its output class distribution.
+In contrast, our method is CLIP-free, can be applied to any pretrained classifier without requiring any annotated data, and explicitly preserves the classifier’s predictive distribution.
+3 Method
+We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design zero-shot concept bottleneck models, which we discuss in Section 3.2.
+3.1 TextUnlock
 A visual classifier assigns an image to a specific category from a predefined set of discrete class labels.
-In ImageNet-1K [6], this set contains $1,000$ class labels.
+For example, in ImageNet-1K (Deng2009ImageNetAL), this set contains $1,000$ class labels.
 Originally, these discrete class labels correspond to class names in text format.
@@
 These classes are typically discretized to facilitate training with cross-entropy.
-However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information.
-Specifically, these embeddings reside in a continuous space where nearby vectors capture related concepts.
-For the “goldfish" example, neighboring vectors might include “freshwater", “fins" and “orange".
-Our method learns to map images into this text embedding space using only class names, thus associating both the class name and its surrounding semantic space with the image.
-This is accomplished through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to output the same distribution scores across all classes, as the original classifier’s class distribution.
-This is done while while keeping both the visual and textual encoders frozen.
+However, when the textual class names are used, they provide an advantage.
+Specifically, when the textual class names are embedded into vector representations (e.g., using a word embedding model or text encoder), they provide semantic information.
+Specifically, These embeddings reside in a continuous space where nearby vectors capture related semantic associations.
+For the “goldfish” example, neighboring vectors might include terms such as “freshwater”, “fins” and “orange”.
+Such associations can be viewed as high-level conceptual attributes that characterize the class.
+Our method learns to map images into this text embedding space using only class names, thus linking both the class name and its surrounding semantic associations with the image.
+This process thus naturally supports unseen words that are not part of the class names (e.g., concepts in the CBM).
+We accomplish this through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to match its distribution with the original classifier’s class distribution.
+This is done while keeping both the visual and textual encoders frozen.
 By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space.
-This allows us to query the visual classifier with any text query beyond the class names.
-We demonstrate the effectiveness of our method with two primary applications.
-1) We build label-free concept bottleneck models for any pretrained classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the newly built linear classifier shares the same reasoning process with the original classifier.
-2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
-In summary, our contributions are as follows: (i) We propose an efficient, label-free method to reformulate any visual classifier such that it can be queried with open-set text queries, maintaining its performance within a negligible margin (average of 0.2 points drop in accuracy) (ii) We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language.
-(iii) On both these applications, our method sets new state-of-the-art results outperforming existing works, including those using CLIP as the visual encoder, while being trained with 400$\times$ less images and 400,000$\times$ less text data.
-2 Related Work
-While CLIP is the predominant approach to interpret vision models through language [2], there exist works that try to decode visual features of models beyond CLIP.
-DeVIL [7] and LIMBER [8] trains an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data.
-Similarly, Natural Language Explanations (NLEs) [9, 10, 11] use annotated textual explanations in place of conventional captions.
-Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process.
-ZS-A2T [12] converts attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective); therefore, it cannot be applied to standalone visual classifiers.
-In contrast, our method decodes visual features of any pretrained classifier in a zero-shot manner without requiring any annotated data, and is explicitly trained to keep the model’s predictive distribution intact.
-Text-to-Concept (T2C) [13] and [14] are the closest related works to ours.
-These works train a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
-Since the linear layer maps features to the CLIP space, these works are strongly biased towards interpreting the CLIP model rather than the original classifier.
-As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image.
-Moreover, these works rely on CLIP supervision during training.
-In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution.
-3 Method
-Figure 2: An overview of our method for reformulating visual classifiers.
-(a) The process of training the MLP mapping between vision and text space.
-A PyTorch-like pseudocode for (a) in included in Section B of the Appendix.
-(b) The process of inference with the adapted visual classifier.
-The text encoder acts as a linear classifier weight generator in our re-formulated classification process.
-Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
+This allows us to query the visual classifier with text queries beyond the class names (e.g., concepts) and obtain concept activations for CBMs.
 Consider an image $I$ and a visual classifier $F$ composed of a visual feature extractor $F_{v}$ and a linear classifier $W$.
@@
 For ImageNet-1K, $K=1000$.
-Consider also any off-the-shelf text encoder $T$ which takes in an input text $l$ and embeds it into a $m-$dimensional vector representation $u\in\mathbb{R}^{m}$.
+Consider also any off-the-shelf text embedding model $T$ which takes in an input text $l$ and embeds it into a $m$-dimensional vector representation $u\in\mathbb{R}^{m}$.
 That is, $u=T(l)$.
 Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
-We propose to learn a simple lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$.
+Figure 1: Overview of our proposed TextUnlock.
+(a) The process of training the MLP mapping between vision and text space with pseudocode given in Appendix Section A.
+(b) The process of inference with the adapted visual classifier.
+The text encoder acts as weight generator for a linear classifier.
+Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
+We propose to learn a lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$.
 That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$.
-Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient.
+Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method parameter-efficient.
 We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}" where {class} represents the class name in text format.
@@
 Here, $U$ acts as weights of the classification layer for our approach.
-We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$.
+We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product.
+between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$.
 Equivalently, this can be performed as a single vector-matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class.
 In other words, $S$ represents the classification logits of our approach.
-The most straightforward way to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution.
-However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions) and sacrificing text interpretability of the original classifier.
+The most straightforward way to training the MLP is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution.
+However, this approach violates two key desiderata: (1) it necessitates annotated data, and (2) re-training the legacy classifier alters its original decision distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions).
 Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored.
-We instead propose to align $S$ to the original distribution $o$ through the cross-entropy loss.
-For a single sample, the loss is given by:
+We instead propose to align $S$ to the original decision distribution $o$ through cross-entropy loss.
+For a single sample, the loss is given by
 $L=-\sum_{i=1}^{K}o_{i}\log\left(\frac{e^{s_{i}}}{\sum_{j=1}^{K}e^{s_{j}}}\right)\;.$ (1)
-This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification.
+This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but instead distill the distribution of the original model to its counterpart vision-language distribution.
 Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters.
@@
 Note also that this loss function naturally encodes the classifier’s relationship across all classes.
-Our training approach is illustrated in Figure 2(a), and we provide a PyTorch-like pseudocode of it in Section B of the Appendix.
-It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section I in the Appendix for more information).
-After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features.
-In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the newly formulated classification task, defined as argmax($\tilde{f}.U^{T}$).
-This is shown in Figure 2(b).
-4 Applications
-In this section, we show how our method can be used to build text-based interpretability applications. At application, all model components (including the MLP) are frozen.
+We provide an illustration of the TexUnlock approach in Figure 1 and a PyTorch-like pseudocode in Section A of the Appendix.
+It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section O in the Appendix for more information).
+After training, the projected visual features and the text encoder features lie in the same space.
+We can therefore query the visual features with any text by finding the alignment score between the embedded text and the projected visual features.
+In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the classification task formulated as argmax($\tilde{f}.U^{T}$), see Figure 1(b).
 3.2 Zero-Shot Concept Bottleneck Models
-Concept Bottleneck Models (CBMs) [1] are a class of inherently interpretable models and have recently attracted significant attention.
-They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions.
+Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed zero-shot CBMs.
+Note that at this stage, all model components (including the MLP) are frozen and no further training is performed.
+CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction.
 In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image.
-In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of opaque features.
-Recently, Label-Free CBMs (LF-CBMs) use CLIP to perform the concept discovery step without annotated image-concept data.
-This is achieved either by using CLIP as a model to provide ground-truth image-concept similarities to train a layer that performs concept discovery [15], or by querying the image features from a set of predefined concepts within the CLIP space [16, 17] and using the cosine similarity scores between the concepts and the image as concept activations.
-However, there are major limitations in this existing line of CBMs.
-Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model.
-Secondly, all CBM approaches to date require training a linear probe on the concept activations, which 1) hinders an on-the-fly ready-to-deploy CBM at test-time directly, and 2) requires training a new $W^{con}$ classifier from scratch for every different set of concepts, hindering flexibility to the choice of concepts.
-Our method solves all the aforementioned issues, formulating zero-shot CBMs for any pretrained classifier.
+In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
+Concept Discovery.
 We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier.
 We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$.
-Following other works [18], and without loss of generality, we use the $Z=20K$ most common words in English [19] as our concept set.
+Following other works (Rao2024DiscoverthenNameTC), and without loss of generality, we use the $Z=20K$ most common words in English (mostcommongoogle) as our concept set.
 These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset.
+We ablate on other concept sets in Appendix Section I for the interested reader.
 To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set.
@@
 In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
+Details of this procedure can be found in Section N of the Appendix.
 This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
+With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot.
+This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
+Figure 2: Building zero-shot CBMs for any pretrained classifier.
+We first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in a zero-shot manner, which results in (c) our final CBM.
+Note that the concept bank only needs to be encoded once.
 We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$.
@@
 We provide an illustration in Figure 2(a).
-Figure 3: (a) Our approach to concept discovery in CBMs.
-(b) Decoding visual features into natural language using any pretrained language decoder (e.g., GPT-2).
-We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features.
-Shown for iteration $j=1$
-We now show how to build the classifier $W^{con}$ that takes in the concept activations and outputs a distribution $S_{cn}$ over classes.
-We build $W^{con}$ in a zero-shot manner.
-Here, zero-shot indicates that no training is required to map concept activations to classes.
+Concept-to-class Prediction.
+The classifier $W^{con}$ takes in concept activations and outputs a distribution $S_{cn}$ over classes.
+We build $W^{con}$ in a zero-shot manner, where here zero-shot indicates that no training is required to map concept activations to classes.
 Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space.
 Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name.
-Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the linear classifier $U$ to obtain the new weights for $W^{con}$.
+Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$.
 That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$.
-Therefore, the weights of $W^{con}$ represent how similar the class name is to all the concepts.
+Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts.
+This process is shown in Figure 2(b).
 In total, the output distribution $S_{cn}$ of the CBM is defined as:
@@
 From Eq.
-2 we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
-Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$.
-Therefore, to convert any classifier to a CBM, all we have to do is plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model.
+2 we make an interesting observation.
+Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
+The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other.
+Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$.
+Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model.
 Eq.
 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts.
-This means our CBMs preserve the basis reasoning process of the original classifier.
-By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner and for any classifier.
-Note that our method is flexible in the concept set.
-The concept set is chosen at test time directly, and the CBM is built directly on-the-fly.
-4.2 Zero-Shot Decoding of Visual Features into Text
-In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into a natural language sentence.
-This offers a text-based interpretation of what the visual features contain.
-We adapt the method introduced in [3] for our purpose.
-Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$.
-That is, $\tilde{f}=\text{MLP}(f)$.
-Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text.
-We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text.
-We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [20] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
-Specifically, we attach a set of learnable “virtual" tokens to $G$.
-Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number.
-$h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$.
-As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens.
-We perform this process for several iterations.
-An overview of this process is shown in Figure 3(b).
-As this process is not the core contribution of our work, we leave more details to Section H of the appendix, which also includes a detailed illustration in Figure 2.
+This means our CBMs preserve the basic reasoning process of the original classifier.
+By this, we obtain zero-shot CBMs that discover concepts and build $W^{con}$ in a zero-shot manner for any classifier (Figure 2(c)).
+An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly.
+This makes our method highly flexible with respect to the chosen concept set.
 4 Experiments
-We first provide results on our newly formulated visual classifiers.
-We use the challenging ImageNet-1K dataset due to the widespread publicly available visual classifiers trained and evaluated on it.
-We apply our method on a diverse set of 40 visual classifiers.
-For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [21], Wide ResNets [22], ResNeXts [23], ShuffleNetv2 [24], EfficientNetv2 [25], Densely Connected Networks (DenseNets) [26], ConvNeXts [27] and ConvNeXtv2 [28].
-For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [29], DINOv2 [30], BeiT [31], the hybrid Convolution-Vision Transformer CvT [32], Swin Transformer [33] and Swin Transformer v2 [34].
-All models are pretrained on ImageNet-1K from the PyTorch [35] and HuggingFace [36] libraries.
-Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K.
-Models with a subscript v2 are trained following the updated PyTorch training recipe [37].
-Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k.
-Both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq.
-1.
-Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K.
-For the text encoder, we use the MiniLM Sentence Encoder [38], with ablation studies on other text encoders presented in Section D of the Appendix.
-Results are presented in Table 1.
-We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column.
-As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models.
-We also perform ablation studies on the MLP to verify its effectiveness in Section C of the Appendix.
-We refer to Appendix Section F for results on other models and to Section G for implementation details.
 Model Top-1 Orig. $\Delta$ Model Top-1 Orig. $\Delta$
-ResNet50 75.80 76.13 --0.33 ResNeXt50-32x4dv2 80.79 80.88 --0.09
+ResNet50 75.80 76.13 -0.33 ResNeXt50-32x4dv2 80.79 80.88 -0.09
 […]
 […]
-ResNet101 77.19 77.37 --0.18 ViT-B/16 80.70 81.07 --0.37
+ResNet101 77.19 77.37 -0.18 ViT-B/16 80.70 81.07 -0.37
 […]
@@
 […]
-DenseNet169 75.46 75.60 --0.14 Swin-Base 83.22 83.58 --0.36
-EfficientNetv2-S 84.04 84.23 --0.19 Swinv2-Base 83.72 84.11 --0.39
-EfficientNetv2-M 84.95 85.11 --0.16 BeiT-B/16 84.54 85.06 --0.52
+DenseNet169 75.46 75.60 -0.14 Swin-Base 83.22 83.58 -0.36
+EfficientNetv2-S 84.04 84.23 -0.19 Swinv2-Base 83.72 84.11 -0.39
+EfficientNetv2-M 84.95 85.11 -0.16 BeiT-B/16 84.54 85.06 -0.52
 […]
-ConvNeXt-Small 83.42 83.62 --0.20 DINOv2-B 84.40 84.22 +0.18
-ConvNeXt-Base 83.88 84.06 --0.18 ConvNeXtV2-B 84.56 84.73 --0.17
+ConvNeXt-Small 83.42 83.62 -0.20 DINOv2-B 84.40 84.22 +0.18
+ConvNeXt-Base 83.88 84.06 -0.18 ConvNeXtV2-B 84.56 84.73 -0.17
 […]
 […]
-Table 1: Performance comparison of our re-formulated classifiers for several models.
+Table 1: TextUnlock-ed visual classifiers maintain classification performance.
+Comparison of our re-formulated classifiers for several models.
 Top-1 indicates our results of the new formulation, and Orig.
 denotes the original Top-1 accuracy.
-$\Delta$ represents the difference between them ($\Delta$ = Top-1 -- Orig).
-Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2.
-As baselines, we compare with 9 methods: LF-CBMs [15], LaBo [16], CDM [17], DCLIP [39], DN-CBMs [18], DCBM [40], Multi-Modal (MM) CBMs [41], CF-CBM [42] and SALF-CBM [43].
-All these methods are supervised and use CLIP-based models either for computing the concept activations or as a backbone.
-Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art.
-Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples.
-The best results are obtained by the ConvNeXtv2 model, which achieve a top-1 accuracy of $86.4$.
-All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
-Figure 4: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction.
-Effectivenss of Concept Interventions: As a supplementary experiment, we also report concept intervention results on our ZS-CBMs in Section E of the Appendix, showing the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
+$\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
+We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock.
+We use the ImageNet-1K benchmark dataset due to the widespread publicly available visual classifiers trained and evaluated on it.
+In Section F of the Appendix, we also report results on other datasets including Places365 (zhou2017places), EuroSAT (helber2019eurosat) and DTD (Cimpoi2013DescribingTI) showing that our method is also applicable to domain-specific and fine-grained datasets.
+We apply TextUnlock on a diverse set of 40 visual classifiers.
+For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He2015DeepRL), Wide ResNets (Zagoruyko2016WRN), ResNeXts (Xie2016AggregatedRT), ShuffleNetv2 (Ma_2018_ECCV), EfficientNetv2 (Tan2021EfficientNetV2SM), Densely Connected Networks (DenseNets) (Huang2016DenselyCC), ConvNeXts (Liu2022ACF) and ConvNeXtv2 (Woo2023ConvNeXtVC).
+For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (dosovitskiy2021an), DINOv2 (oquab2024dinov), BeiT (bao2022beit), the hybrid Convolution-Vision Transformer CvT (Wu2021CvTIC), Swin Transformer (Liu2021SwinTH) and Swin Transformer v2 (Liu2021SwinTV).
+All models are pretrained on ImageNet-1K from the PyTorch (torchvision2016) and HuggingFace (wolf-etal-2020-transformers) libraries.
+Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K.
+Models with a subscript v2 are trained following the updated PyTorch training recipe (PytorchNewTrainRecepie).
+Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k.
+When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq.
+1.
+Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K.
+For the text encoder, we use the MiniLM Sentence Encoder (wang2020minilm) as it is powerful, fast and efficient.
+We provide ablation studies on other text encoders in Section D of the Appendix.
+Results are presented in Table 1.
+We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column.
+As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models.
+We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section C of the Appendix (see Section C.1 for the role and impact of the MLP and Section C.2 for its design hyperparameters).
+We also evaluate our transformed classifier robustness to prompt variations in Section H of the Appendix.
+For results on other models, we refer to Section L of the Appendix, and for implementation details, to Section J of the Appendix.
+[…]
+Effectivenss of Concept Interventions.
+Another common way to evaluate interpretability of CBMs is concept intervention.
+We report concept intervention results on our ZS-CBMs in Section E of the Appendix.
+In this way, we show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
 Supervised CBMs
 Method Model Top-1 Method Model Top-1
-LF-CBM CLIP ResNet50 67.5 DN-CBM CLIP ViT-B/16 79.5
-LF-CBM CLIP ViT-B/16 75.4 SALF-CBM-S CLIP ViT-B/16 75.3
-LaBo CLIP ResNet50 68.9 SALF-CBM-D CLIP ViT-B/16 76.3
-LaBo CLIP ViT-B/16 78.9 DCBM-SAM2 CLIP ViT-L/14 77.9
-CDM CLIP ResNet50 72.2 DCBM-GDINO CLIP ViT-L/14 77.4
-CDM CLIP ViT-B/16 79.3 DCBM-RCNN CLIP ViT-L/14 77.8
-DCLIP CLIP ViT-B/32 63.0 MM-CBM CLIP ViT-B/16 77.9
-DCLIP CLIP ViT-B/16 68.0 MM-ProtoSim CLIP ViT-B/16 78.8
-DCLIP CLIP ViT-L/14 75.0 CF-CBM-H CLIP ViT-B/16 77.4
-DN-CBM CLIP ResNet50 72.9 CF-CBM-L CLIP ViT-B/16 78.4
+LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
+LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
+CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
+DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
+DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
+DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
 Zero-Shot CBMs (Ours)
@@
 […]
-Table 2: Performance of Supervised and Zero-Shot CBMs on the ImageNet validation set.
-In Figure 4, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance.
+Table 2: Our zero-shot CBMs outperform CLIP-based counterparts.
+Accuracy of Supervised and Zero-Shot CBMs (ZS-CBMs, ours) on ImageNet validation set.
+Similar backbones are color-coded.
+Best within backbone family is underlined, overall best is bold.
+In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis.
 The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class.
 We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder.
-All examples use the LF-CBM concept set [15].
+All examples use the LF-CBM concept set (oikarinen2023labelfree).
 By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab.
-Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [44] regarding the “dumbbell” class.
+Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature (Samek2019TowardsEA) regarding the “dumbbell” class.
 Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it.
 With our method, we can obtain a textual interpretation of the biases that the original classifier learns.
-We also provide qualitative examples of global class-wise concepts detected in Section K of the appendix.
-Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features.
-Since the ground-truth content of visual features is unknown—that is, we do not know exactly what the model uses for reasoning—we rely on COCO captions [45] for evaluation.
-While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose, and represent the best available alternative for evaluation.
-Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person).
-Therefore, it also serves as a way to evaluate generalization of our method to other datasets, given that we only used the ImageNet class names for training.
-Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark.
-We present results on the widely used “Karpathy test split" with various vision classifiers.
-As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap [3] and ConZIC [46], both which use CLIP.
-For evaluation, we employ standard natural language generation metrics: BLEU-4 (B@4) [47], METEOR (M) [48], ROUGE-L (R-L) [49], CIDEr (C) [50], and SPICE (S) [51].
-Results are shown in Table 3(a).
-ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems.
-Even with a simple ResNet-50 vision encoder trained on ImageNet-1K (1.2 million images), our approach outperforms the baseline methods on CIDEr and SPICE, despite the latter utilizing the significantly more powerful CLIP vision encoder, trained on 400 million image-text pairs.
-We also present qualitative examples of the decoded visual features for different models in Figure 5.
-From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables.
-In contrast, ViT-B/16 and DINOv2 focus exclusively on the dog and its characteristics.
-Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason.
-This application allows human-friendly interpretations of visual features, understandable even to a layman user.
+We also provide qualitative examples of global class-wise concepts detected in Section P of the appendix.
+Figure 3: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the on the x-axis.
+5 Zero-Shot Image Captioning
+We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models.
+Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space.
+Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier.
+We adapt the method introduced in ZeroCap (Tewel2021ZeroCapZI) for our purpose.
+Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features.
+We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$.
+That is, $\tilde{f}=\text{MLP}(f)$.
+Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text.
+We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text.
+We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li2021PrefixTuningOC) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
+Specifically, we attach a set of learnable “virtual" tokens to $G$.
+Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number.
+$h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$.
+As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens.
+We perform this process for several iterations.
+As this process is not the core contribution of our work, we leave more details to Section M of the appendix, which also includes a detailed illustration in Figure 2.
+[…]
+Figure 4: Qualitative examples of zero-shot image captioning.
 Model B4 M R-L C S
@@
 […]
-(a) Zero-Shot Image Captioning Performance
-Model B4 M R-L C S
-ZeroCap 2.6 11.5 14.6 5.5
-ConZIC 1.3 11.5 12.8 5.2
-Ours
-DenseNet161 4.20 12.5 30.1 17.0 6.6
-ResNet50 4.10 12.5 30.1 17.0 6.6
 […]
-[…]
-ResNet101v2 4.30 12.6 30.1 18.0 6.8
-[…]
-[…]
-EfficientNetv2-S 4.40 12.7 30.4 18.6 6.9
-[…]
-[…]
-BeiT-B/16 4.50 12.8 30.3 18.9 7.1
-DINOv2-Base 4.60 13.0 30.7 18.7 7.1
-(b) Zero-Shot Compositional Captioning Performance
-Table 3: Comparison of Zero-Shot Captioning Performance
-Figure 5: Qualitative examples of decoded visual features from different visual classifiers, into text.
-Note that our results in Table 3(a) are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics.
+Table 3: Zero-Shot Image Captioning Performance
+Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics.
 However, it is important to note that B4 and M are n-gram overlap-based metrics.
 They assume that the generated caption follows a specific structure and style.
-We verify this by revisiting an old image captioning paradigm termed as compositional captioning [52], and later revived with deep learning methods [53].
-In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence.
-With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner.
-Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3(a), and feed them, along with their similarity scores, to an LLM.
-For the concepts, we use the same concept set as in Section 4.1.
-We also add a list of the most common verbs [54] in English to the pool.
-This allows us to cover all possible words and interactions.
-We prompt the LLM to utilize the provided information to compose a sentence, given a very few in-context examples from the COCO captioning training set (in our experiments, we use 9 examples).
-This allows us to generate sentences adhering to a specific style and structure.
-For this experiment, we used GPT4o-mini [55] as our LLM, as it is fast and cost-efficient.
-In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances.
-We provide details on the exact prompt we used in Appendix Section J.
-Results are shown in Table 3(b).
-While results on CiDEr and SPICE are incremental compared to results in Table 3(a), the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3(a) compared to baseline methods.
-Also note that, since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains (see Section J.1 in the appendix).
+We verify this hypothesis by applying compositional image captioning (Kulkarni2011BabyTU; Lu2018NeuralBT) with in-context learning.
+Specifically, a set of image-grounded verbs and concepts (e.g., attributes, objects) are first detected using the concept discovery step described in Section 3.2, and they are fed along with six in-context examples from the COCO captioning set to a language model, capable of imitating the style and structure of the in-context examples, in order to compose them into a natural sounding sentence.
+This also demonstrates the direct application of our approach to compositional image captioning.
+The last row of Table 3 with the superscript com shows that the (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M compared to baseline methods.
+We provide more details and results on other models on this paradigm of compositional image captioning in Section K of the Appendix.
+We further show how we can use this approach to generate captions tailored to any domain (see Section K.1 in the Appendix).
 6 Conclusion
-We presented a method to transform visual classifiers such that they can be queried via open-set text queries.
-We achieve state-of-the-art results in two applications, i.e., zero-shot CBMs and decoding of visual features into natural language sentences.
-Our method can be applied to various other applications that were previously restricted to CLIP models.
-Our work removes this restriction, being applicable to any visual classifier.
-Finally, as with any research work, this study has its own limitations which are discussed in Appendix Section A.
-Appendix A Limitations
+We introduced a method for transforming any frozen visual classification model into a Concept Bottleneck Model (CBM).
+We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart.
+This method further enables a CLIP-free, label-free conversion of the frozen visual classification model into a CBM, and further allows us to derive zero-shot concept bottleneck concept-to-class classifiers (without training the linear classifier), Applied on 40 different models, our method outperforms supervised CLIP-based CBMs and achieves state-of-the-art results.
+TextUnlock further allows us to perform zero-shot image captioning on any visual classification model, also achieving state-of-the-art results and surpassing zero-shot captioning baselines based on CLIP.
+Finally, as with any research work, this study has its own limitations which are discussed in Section B of the Appendix.
+Appendix A Pseudocode of TextUnlock
+We provide a PyTorch-like pseudocode of TextUnlock in Listing 1.
+# text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
+# classifier: linear classifier weights of a frozen vision_encoder, shape (visual_dim, num_classes)
+# mlp: trainable MLP from visual_dim -> text_dim
+# images: batch of B images, shape (N, 3, height, width)
+visual_feats = vision_encoder(images) # (N, visual_dim)
+logits = visual_feats @ classifier # (N, num_classes)
+original_dist = softmax(logits, dim=-1) # (N, num_classes)
+mapped_feats = mlp(visual_feats) # (N, text_dim)
+mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
+text_feats = l2_norm(text_feats) # (N, text_dim)
+pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
+pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
+# cross entropy with original model’s soft distribution
+loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
+loss.backward() # only mapper parameters are updated
+Listing 1: PyTorch-like pseudocode for TextUnlock
+Appendix B Limitations
 As with any research work, our study has its own limitations that should be transparent and acknowledged.
-In particular, we identify two primary limitations of our method: the first concerns wrong semantic associations of class names in CBMs, while the second pertains to the limited generalization of our method to fine-grained datasets.
-We start by addressing the first limitation of wrong semantic associations of class names in CBMs.
+In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs.
 This issue is closely tied to the choice of concept set used.
@@
     This issue also appears in CLIP-based CBMs and hence not unique to our approach.
-Figure 1: Limitations of our method is wrong semantic concept association in CBMs
-The second limitation is the lack of generalization to fine-grained datasets.
-We test our ImageNet-trained reformulated classifiers to perform zero-shot transfer to the Places365 dataset [56] for scene classification.
-It is important to note that Places365 is known to induce low performance, even when using supervised training with the powerful CLIP model, making it a challenging classification dataset.
-To perform zero-shot transfer, we convert the class names of Places365 into text prompts of the form: “a scene of a {class} location" which are then encoded using the text encoder $T$ to obtain the classification weights $U_{places}$.
-The image is encoded with the respective vision encoder to obtain the visual features $f$, which are then projected via the MLP to yield $\tilde{f}$.
-Zero-shot classification is performed using $\text{argmax}(\tilde{f}.U^{T}_{places})$.
-Results on the Places365 validation set are shown in Table 1.
-We report the supervised learning performance by training a linear probe on top of the CLIP model features.
-The supervised top-1 accuracy is 53.40 and 55.10 for ResNet50 and ViT-B/16 CLIP models, respectively.
-As demonstrated, these low top-1 accuracies suggest that Places365 is a particularly challenging dataset for classification.
-We then report the zero-shot performance of our method on several visual classifiers.
-The best-performing model achieves a top-1 zero-shot accuracy of 14.36%, much lower than the supervised performance.
-However, we emphasize that the observed low generalization performance is not attributed to our formulation; rather, it stems from the inherent limitations of the classifier itself.
-If the classifier (e.g., ResNet101) fails to generalize to out-of-distribution (OOD) fine-grained classes—as is the case—then it is unreasonable to expect the text-transformed classifier to perform any better.
-Since our transformed classifier is a replication of the original classifier, it inherits all of its limitations.
-In fact, expecting improved performance from the transformed classifier would be counterintuitive, as it would indicate a departure from the original classifier’s underlying reasoning.
-Moreover, as demonstrated in Table 1, models that benefit from large-scale pretraining tend to achieve superior OOD fine-grained performance, a trend that is consistent with findings in the OOD literature [57].
-Finally, our approach performs significantly better than a randomly shuffled classifier where we shuffle its weights across classes.
-Model Top-1 (%) Top-5 (%)
-Supervised
-CLIP RN50 Linear Probe 53.40 -
-CLIP ViT-B Linear Probe 55.10 -
-Zero-Shot
-Shuffled Classifier 0.24 1.54
-DenseNet161 10.55 25.18
-ResNet50 10.58 25.14
-ResNet101 10.97 25.42
-ViT-B/16 11.47 26.53
-ResNet50v2 11.66 27.40
-ConvNeXt-Base 11.76 27.12
-Swin-Base 12.12 26.83
-Swinv2-Base 12.13 27.18
-EfficientNetv2-M 12.31 28.33
-ViT-B/16v2 13.11 29.54
-DINOv2-Base 13.96 32.35
-ConvNeXtv2 (pt) 14.05 30.92
-BeiT-L/16 14.15 31.08
-ViT-B/16 (pt) 14.22 32.18
-ConvNeXt (pt) 14.36 31.64
-Table 1: Top-1 and Top-5 zero-shot transfer results to Places365 for various models
-Appendix B Pseudocode of our method
-We provide a PyTorch-like pseudocode of our training approach in Listing 1.
-# text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
-# classifier: linear classifier weights of a frozen vision_encoder, shape (visual_dim, num_classes)
-# mlp: trainable MLP from visual_dim -> text_dim
-# images: batch of B images, shape (N, 3, height, width)
-visual_feats = vision_encoder(images) # (N, visual_dim)
-logits = visual_feats @ classifier # (N, num_classes)
-original_dist = softmax(logits, dim=-1) # (N, num_classes)
-mapped_feats = mlp(visual_feats) # (N, text_dim)
-mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
-text_feats = l2_norm(text_feats) # (N, text_dim)
-pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
-pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
-# cross entropy with original model’s soft distribution
-loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
-loss.backward() # only mapper parameters are updated
-Listing 1: PyTorch-like pseudocode for our method
+Figure 1: Limitations of our method in wrong semantic concept association
 Appendix C Ablation Studies of the MLP
-In this section, we evaluate the impact of the MLP and verify its effectiveness.
+C.1 Role and Impact of the MLP
+We first evaluate the role and impact of the MLP to verify its contribution and that it learns meaningful transformations, and does not collapse into a trivial function.
 We perform the following ablation studies: 1) Mean Ablation: For an image, we replace the input features to the MLP with a constant mean of the features calculated across the full ImageNet validation set.
@@
 Table 1: Ablation studies of the MLP.
+C.2 MLP Design
+Next, we perform an ablation study over the number of layers and output dimension scaling factor (dim_out_factor) of the MLP.
+dim_out_factor specifies how much to scale the previous layer’s dimensionality.
+For example, in a ViT-B/16 model the hidden dimension is 768; setting dim_out_factor = 2 therefore expands the projection from 768 to 2 $\times$ 768 = 1536.
+Results are shown in Table 2 for a ResNet50 model.
+We observe that the 2 layers and a output dimension scaling factor of 2 provides the best results.
+Layers dim_out_factor Top-1 (%)
+1 1 72.48
+1 2 74.01
+2 1 75.41
+2 2 75.80
+Table 2: Performance comparison of the MLP for different layer and dimension configurations.
 Appendix D Ablation Studies on the Text Encoder
-In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library [58] with the ResNet50 visual classifier.
+In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library (reimers-2019-sentence-bert) with the ResNet50 visual classifier.
 We observe that the choice of the text encoder has minimal effect on the performance.
@@
 Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
-The Waterbirds-100 dataset [59] is a standard dataset used in previous works [18] to conduct CBM intervention experiments.
+The Waterbirds-100 dataset (Sagawa2020Distributionally) is a standard dataset used in previous works (Rao2024DiscoverthenNameTC) to conduct CBM intervention experiments.
 It is a binary classification dataset of two classes: waterbirds and landbirds.
@@
 By building a CBM, we can correct this bias by intervening in concepts in the CB layer.
-However, there are two challenges associated with using the Waterbirds dataset in our work, and therefore we curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set.
-We provide more details about this process in Section L.
+We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set.
 Our curated dataset includes 140 validation images (70 for each class).
-We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from [18] which includes a collection of bird-related concepts and a collection of land-related concepts.
+We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from (Rao2024DiscoverthenNameTC) which includes a collection of bird-related concepts and a collection of land-related concepts.
 The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification.
-To correct this, we intervene in the concepts in the CB layer, following the setup from [18].
+To correct this, we intervene in the concepts in the CB layer, following the setup from (Rao2024DiscoverthenNameTC).
 Intervention R (Int.
@@
 We also conduct concept intervention experiments with a more challenging multi-class setup.
-Unlike the Waterbirds dataset where bias-correlation issues are assumed, we make no such assumption here.
+Unlike the Waterbirds dataset where bias‑correlation issues are assumed, we make no such assumption here.
 We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy.
 A drop in accuracy indicates that these concepts are important for prediction.
-We select a subset of 10 classes from ImageNet following [60].
-We use the original ImageNet validation images for those classes, rather than the validation set from [60], because the latter contains original ImageNet training examples that our model has already seen.
+We select a subset of 10 classes from ImageNet following (Howard_Imagenette_2019).
+We use the original ImageNet validation images for those classes, rather than the validation set from (Howard_Imagenette_2019), because the latter contains original ImageNet training examples that our model has already seen.
 We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts.
-The classes and concepts we used are provided in Section M of the supplementary material.
+The classes and concepts we used are provided in Section Q of the supplementary material.
 We then measure CBM accuracy before and after intervention.
@@
 As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
-Model Orig. CBM Int. (R)$\downarrow$ Int. (K)$\uparrow$
+[…]
 BeiT-B/16 54.29 41.43 (-12.86) 58.57 (+4.28)
@@
 Table 4: CBM Interventions on the standard Waterbirds dataset
-Model Orig. CBM Int. (R)$\downarrow$
+[…]
 ResNet50 96.80 76.00 (-20.8)
@@
 Table 5: CBM Interventions in a multi-class setup
-Appendix F Performance on Additional Models
-We report performance on additional models that were not included in the main manuscript in Table 6.
-Model Top-1 Orig. $\Delta$
-ConvNeXt-Tiny 82.19 82.52 --0.33
-ViT-B/32 75.40 75.91 --0.51
-Swin-Small 82.63 83.20 --0.57
-Swinv2-Tiny 81.44 82.07 --0.63
-CvT-21 80.45 81.27 --0.82
-Swinv2-Small 83.32 83.71 --0.39
+Appendix F Transformed Classifier Results on Other Datasets
+We conduct extra experiments on 3 datasets.
+Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images).
+For each dataset, we show the top-1 accuracy of the classifier’s original performance and the top-1 accuracy of our transformed classifier.
+Results are shown in Table 6.
+Dataset Model Original (%) Transformed (%)
+Places365 ResNet50 54.77 53.90
+DenseNet161 56.13 55.54
+EuroSAT ResNet50 93.95 94.23
+WideResNet101 93.78 93.88
+ViT-B/16 93.67 93.53
+DTD ResNet50 69.47 69.26
+WideResNet101 68.62 68.14
+ViT-B/16 69.95 69.68
+Table 6: Results of our transformed classifier on Places365, EuroSAT, and DTD datasets
+Appendix G CBM Results on Other Datasets
+We report CBM results on additional datasets: Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images).
+When a baseline method is also reported on that dataset, we include it as a baseline.
+Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline.
+All baselines use the same concept set.
+Results are shown in Table 7.
+For Places 365, we can see that the transformed DenseNet161 classifier trained only on ImageNet, outperforms supervised CLIP-based ResNet and ViTs CBM methods in a zero-shot manner.
+The same applies for EuroSAT and DTD with our transformed Image-Net only trained classifiers.
+Therefore, the experiments show that our method scales to domain-specific and fine-grained datasets.
+Dataset Method Model Acc (%)
+Places365 Supervised
+DCLIP CLIP-ResNet50 37.90
+DCLIP CLIP-ViT-B/16 40.30
+LF-CBM CLIP-ResNet50 49.00
+LF-CBM CLIP-ViT-B/16 50.60
+CDM CLIP-ResNet50 52.70
+CDM CLIP-ViT-B/16 52.60
+Zero-Shot
+Ours ResNet50 51.57
+Ours DenseNet161 53.42
+EuroSAT Supervised
+Baseline CLIP-ResNet50 86.27
+Baseline CLIP-ViT-B/16 88.57
+Zero-Shot
+Ours ResNet50 94.22
+Ours WideResNet101 94.12
+Ours ViT-B/16 93.65
+DTD Supervised
+Baseline CLIP-ResNet50 57.77
+Baseline CLIP-ViT-B/16 61.86
+Zero-Shot
+Ours ResNet50 68.88
+Ours WideResNet101 66.97
+Ours ViT-B/16 68.46
+Table 7: CBM results on Places365, EuroSAT, and DTD datasets
+Appendix H Prompt Variations
+We evaluate the robustness of our transformed classifier to variations in text prompts.
+Using the transformed ViT-B/16, we provide a diverse set of prompts and measure the Top-1 accuracy on ImageNet.
+Results are shown in Table 8.
+We order the prompts from highest to lowest scoring.
+We find that our transformed classifier is robust to text prompts.
+The worst prompt degrades the baseline accuracy by only 0.36 points
+Prompt Top-1 Accuracy (%)
+an image of a {} 80.70
+a photo of one {}. 80.67
+a photo of a {}. 80.66
+a close-up photo of a {}. 80.63
+a black and white photo of a {} 80.63
+a close-up photo of the {}. 80.62
+a cropped photo of a {}. 80.61
+a good photo of a {}. 80.61
+a dark photo of a {}. 80.61
+a bright photo of a {}. 80.60
+a bright photo of the {}. 80.59
+a bad photo of a {}. 80.57
+a blurry photo of a {}. 80.57
+a pixelated photo of a {}. 80.55
+a photo of many {}. 80.54
+a low-resolution photo of the {} 80.54
+a low-resolution photo of a {}. 80.52
+a photo of my {}. 80.45
+a jpeg-corrupted photo of a {}. 80.34
+Table 8: Effect of different text prompts on Top-1 accuracy using ViT-B/16.
+Appendix I Other Concept Sets
+Early works such as label-free CBMs use LLMs to automatically generate a set of concepts for a target class.
+In recent works such as DN-CBM (Rao2024DiscoverthenNameTC) and DCBM (Knab2024DCBMDV), it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations.
+In Table 9, we compare general pre-defined and LLM-generated concept sets.
+Across different types of architectures, general, pre-defined concept sets outperform the LLM-generated ones by several percent.
+Please note that while earlier works such as LF-CBMs employ LLM-generated concept sets (which underperform compared to the general, pre-defined concept set), all results reported in our main manuscript—including those from prior works—are presented using the general, pre-defined concept set to ensure a fair comparison.
+Model Predefined LLM-generated
+Swinv2-Base 82.60 80.34
+ConvNext-Base 82.80 80.24
 […]
-Table 6: Performance of our reformulated classifiers for additional models
-Appendix G Implementation Details
-For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library [58].
+ViT-B/16 79.30 76.37
+Table 9: Comparison of Predefined and LLM-generated concept sets on CBMs
+Appendix J Implementation Details
+For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (reimers-2019-sentence-bert).
 This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs.
 It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
-Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization [61], a GELU activation function [62] and Dropout [63] with a drop probability of 0.5.
+Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba2016LayerN), a GELU activation function (Hendrycks2016GaussianEL) and Dropout (Srivastava2014DropoutAS) with a drop probability of 0.5.
 The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function.
 The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder).
-We train the MLP projector with a batch size of 256 using the ADAM optimizer [64] with a learning rate of 1e-4 that decays using a cosine schedule [65] over the total number of epochs.
+We train the MLP projector with a batch size of 256 using the ADAM optimizer (KingBa15) with a learning rate of 1e-4 that decays using a cosine schedule (loshchilov2017sgdr) over the total number of epochs.
 We follow the original image sizes that the classifier was trained on.
@@
 Each model is trained on a single NVIDIA GeForce RTX 2080 Ti GPU.
-Appendix H Process of Decoding Visual Features
-We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$.
-A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$.
-To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning [20], which prepends learnable tokens in the embedding space.
-We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly.
-Our method builds upon the work of [3].
-A high-level overview of this process is illustrated in Figure 2.
-Using a pre-trained language model $G$, we prepend randomly initialized learnable tokens, referred to as prefixes, which guide $G$ to produce text that maximizes alignment with visual features.
-These learnable prefixes function as key-value pairs in each attention block, ensuring that every generated word can attend to them.
-For each iteration $j$, at a timestep $ts$, we sample the top-$Q$ tokens from the output distribution of $G$, denoted as $G_{out}$, which serve as possible continuations for the sentence.
-These $Q$ candidate sentences are then encoded by a text encoder $T$, mapping them into the same embedding space as $\tilde{f}$.
-We compute the cosine similarity between each encoded sentence and $\tilde{f}$, resulting in $Q$ similarity scores.
-These scores are normalized with softmax and define a target distribution used to train $G_{out}$ via Cross-Entropy loss.
-The learnable prefixes are updated through backpropagation.
-With the updated prefixes, $G$ is run again, and the most probable token is selected as the next word.
-This process is repeated for a predefined number of timesteps (up to the desired sentence length) or until the <.><.>< .
-> token is generated.
-At the end of each iteration, a full sentence is generated.
-We conduct this process for 20 iterations, generating 20 sentences in total.
-The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
-We also add the fluency loss from [3] as well as other token processing operations.
-We refer readers to [3] for more information.
-We use the small GPT-2 of 124M parameters as $G$.
-We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
-Figure 2: The process used to decode visual features of an image. The process is shown for the first timestep $ts=1$ with a hard prompt set as “an image of a".
-Appendix I Using only class names
-We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP.
-In practice, we can go beyond class names by using resources like a class hierarchy from WordNet [66] (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL [67] or VCVD[39].
-However, this approach would be considered as “cheating.
-The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes.
-Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn.
-Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation.
-For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts.
-For these reasons, we refrain from using any other additional information than the class names.
-We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
-Appendix J Compositional Captioning
-The prompt we used for compositional captioning is as follows:
+Appendix K Compositional Image Captioning
+Compositional Image Captioning is an old image captioning paradigm termed (Kulkarni2011BabyTU), later revived with deep learning methods (Lu2018NeuralBT).
+In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence.
+With the current advancements of Large Language Models (LLMs) and their powerful capabilities, we use an LLM as a compositioner.
+Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 3.2 and shown in Figure 2(a), and feed them, along with their similarity scores, to an LLM.
+For the concepts, we use the same concept set as in Section 3.2.
+We also add a list of the most common verbs (commonverbs) in English to the pool.
+This allows us to cover all possible words and interactions.
+We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 9 examples).
+This allows us to generate sentences adhering to a specific style and structure.
+For this experiment, we used GPT4o-mini (gpt4omini) as our LLM, as it is fast and cost-efficient.
+In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importance.
+We use the following prompt:
 I will give you several attributes and verbs that are included in an image, each with a score.
@@
 Here are the attributes and scores: {detected concepts with scores}, and these are the verbs and scores: {detected verbs with scores}.
-J.1 Domain-Specific Compositional Captioning
-Next, we explore alternative concept sets in Compositional Captioning.
+Results are shown in Table 10.
+Model B4 M R-L C S
+ZeroCap 2.6 11.5 14.6 5.5
+ConZIC 1.3 11.5 12.8 5.2
+Ours
+DenseNet161 4.20 12.5 30.1 17.0 6.6
+ResNet50 4.10 12.5 30.1 17.0 6.6
+[…]
+[…]
+ResNet101v2 4.30 12.6 30.1 18.0 6.8
+[…]
+[…]
+EfficientNetv2-S 4.40 12.7 30.4 18.6 6.9
+[…]
+[…]
+BeiT-B/16 4.50 12.8 30.3 18.9 7.1
+DINOv2-Base 4.60 13.0 30.7 18.7 7.1
+Table 10: Zero-Shot Compositional Captioning Performance
+While the results on CiDEr and SPICE are incremental compared to the results in Table 3, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis that the low scores of B4 and M were attributed to the specific caption style and structure of the respective dataset (COCO).
+K.1 Domain-Specific Compositional Captioning
+Since the LLM we used for compositional image captioning is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains.
+We explore alternative concept sets in Compositional Captioning.
 In the main manuscript, we reported results using the 20,000 most common English words as our concept set.
 Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative.
-This allows for the generation of captions (here, decoded visual features) tailored to a specific domain.
+This allows for the generation of captions tailored to a specific domain.
 Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset.
 Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset.
-We use the ImageNet-specific concept set from [15] and report zero-shot captioning performance in Table 7.
+We use the ImageNet-specific concept set from (oikarinen2023labelfree) and report zero-shot captioning performance in Table 11.
 As shown, we observe a decrease in all metrics.
-This shows that our method can readily decode visual features into text for any domain.
+This shows that our method can readily produce captions for any domain.
 Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
@@
 Table 11: Composition Captioning Performance using the ImageNet-specific LF-CBM concept set
-Appendix K Global Class-Wise Qualitative Examples
+Appendix L Performance on Additional Models
+We report performance on additional models that were not included in the main manuscript in Table 12.
+Model Top-1 Orig. $\Delta$
+ConvNeXt-Tiny 82.19 82.52 -0.33
+ViT-B/32 75.40 75.91 -0.51
+Swin-Small 82.63 83.20 -0.57
+Swinv2-Tiny 81.44 82.07 -0.63
+CvT-21 80.45 81.27 -0.82
+Swinv2-Small 83.32 83.71 -0.39
+[…]
+Table 12: Performance of our reformulated classifiers for additional models
+Appendix M Process of Zero-Shot Image Captioning
+We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$.
+A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$.
+To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li2021PrefixTuningOC), which prepends learnable tokens in the embedding space.
+We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly.
+Our method builds upon the work of (Tewel2021ZeroCapZI).
+A high-level overview of this process is illustrated in Figure 2.
+Using a pre-trained language model $G$, we prepend randomly initialized learnable tokens, referred to as prefixes, which guide $G$ to produce text that maximizes alignment with visual features.
+These learnable prefixes function as key-value pairs in each attention block, ensuring that every generated word can attend to them.
+For each iteration $j$, at a timestep $ts$, we sample the top-$Q$ tokens from the output distribution of $G$, denoted as $G_{out}$, which serve as possible continuations for the sentence.
+These $Q$ candidate sentences are then encoded by a text encoder $T$, mapping them into the same embedding space as $\tilde{f}$.
+We compute the cosine similarity between each encoded sentence and $\tilde{f}$, resulting in $Q$ similarity scores.
+These scores are normalized with softmax and define a target distribution used to train $G_{out}$ via Cross-Entropy loss.
+The learnable prefixes are updated through backpropagation.
+With the updated prefixes, $G$ is run again, and the most probable token is selected as the next word.
+This process is repeated for a predefined number of timesteps (up to the desired sentence length) or until the <.><.> token is generated.
+At the end of each iteration, a full sentence is generated.
+We conduct this process for 20 iterations, generating 20 sentences in total.
+The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
+We also add the fluency loss from (Tewel2021ZeroCapZI) as well as other token processing operations.
+We refer readers to (Tewel2021ZeroCapZI) for more information.
+We use the small GPT-2 of 124M parameters as $G$.
+We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
+Figure 2: The process used to generate zero-shot captions using any pretrained language decoder (e.g., GPT-2).
+The process is shown for the first timestep ($ts=1$) and first iteration ($j=1$) with a hard prompt set as “an image of a".
+We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features.
+Appendix N Concept Filtering
+One of our concept filtering procedures requires us to know the terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
+In order to obtain this information, we used an LLM (gpt-4o-mini) with the following prompt:
+Provide your answer to below as single-word comma separated.
+If you need to provide a term composed of 2 words, then separate each into a single word.
+For each class, provide its synonyms and closely related names (e.g., other species of the category, its superclasses such as bird, fish, dog, cat, animal...etc).
+Here is an example.
+The ImageNet class is: tench, tinca tinca.
+Answer: fish,animal,cyprinid,carp,vertebrate.
+Appendix O Using only class names
+We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP.
+In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin1998WordNetAE) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt2022WhatDA) or VCVD(Menon2022VisualCV).
+However, this approach would be considered as “cheating.
+The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes.
+Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn.
+Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation.
+For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts.
+For these reasons, we refrain from using any other additional information than the class names.
+We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
+Appendix P Global Class-Wise Qualitative Examples
 In Figure 3, we present a probability distribution of global class-wise concepts.
@@
 We highlight in yellow the top concepts in "hammerhead shark" that are not present in "tiger shark", and therefore distinctive to "hammerhead shark".
-Appendix L Waterbirds/Landbirds Dataset
-There are two challenges associated with using the Waterbirds-100 dataset [59] in our work.
-First, our method cannot transform a classifier trained on this dataset because learning meaningful text representations from merely two labels (waterbird and landbird) with our MLP is not feasible.
-Second, we cannot evaluate our ImageNet-trained models on this dataset due to its highly artificial nature, which results in numerous severe out-of-distribution samples for classifiers trained on ImageNet.
-Therefore, we manually curated our own waterbird/landbird dataset using ImageNet validation images.
-Specifically, for the waterbird images, we consider classes of birds from ImageNet that are at least 90% of the time found in water backgrounds— a statistic verified by manually inspecting 100 random training images of those birds.
-For those birds, we then select their images in the ImageNet validation set that appear on land backgrounds.
-We perform a similar procedure for the landbird class.
-For the landbird class, we encountered a problem.
-Although many ImageNet bird classes nearly always feature land backgrounds in their training images, we were unable to find a substantial number of validation images depicting these birds against water backgrounds (e.g., we could not find any image of the bird robin with water background in the ImageNet validation set).
-In order to solve this issue, we utilized the Stable Diffusion 2.1 [68] text-to-image generative model to generate images of those birds on water backgrounds.
-We ensured that the generated images of those birds have the correct physical and distinctive features of the bird.
-For all images, we always ensure that the background (water/land) is clearly visible.
-This leads us to a validation dataset of 140 images (70 images for each class).
-Specifically, for the waterbird class, all 70 images coming from the ImageNet validation set.
-For the landbirds class, 15 images come from the ImageNet validation set, and 55 are generated with Stable Diffusion.
-Appendix M Multi-class CBM Intervention Classes and Concepts
+Appendix Q Multi-class CBM Intervention Classes and Concepts
 Class Concepts

### v4

@@
-CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models
+CLIP-Free, Label Free, Unsupervised Concept Bottleneck Models
 Abstract
-Concept Bottleneck Models (CBMs) map dense, high-dimensional feature representations into a set of human-interpretable concepts which are then combined linearly to make a prediction.
-However, modern CBMs rely on the CLIP model to establish a mapping from dense feature representations to textual concepts, and it remains unclear how to design CBMs for models other than CLIP.
+Concept Bottleneck Models (CBMs) map dense feature representations into human-interpretable concepts which are then combined linearly to make a prediction.
+However, modern CBMs rely on the CLIP model to obtain image-concept annotations, and it remains unclear how to design CBMs without the CLIP bottleneck.
 Methods that do not use CLIP instead require manual, labor intensive annotation to associate feature representations with concepts.
 Furthermore, all CBMs necessitate training a linear classifier to map the extracted concepts to class labels.
-In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in a zero-shot manner.
+In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in an unsupervised manner.
 Our method is formulated by aligning the original classifier’s distribution (over discrete class indices) with its corresponding vision-language counterpart distribution derived from textual class names, while preserving the classifier’s performance.
 The approach requires no ground-truth image–class annotations, and is highly data-efficient and preserves the classifiers reasoning process.
-Applied and tested on over 40 visual classifiers, our resulting CLIP-free, zero-shot CBM sets a new state of the art, surpassing even supervised CLIP-based CBMs.
-Finally, we also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art results.
+Applied and tested on over 40 visual classifiers, our resulting unsupervised, label-free and CLIP-free CBM (U-F2-CBM) sets a new state of the art, surpassing even supervised CLIP-based CBMs.
+We also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art.
 1 Introduction
 Visual classifiers predict a class as a linear combination of dense, high‑dimensional visual feature vectors that are difficult to interpret by humans.
-Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation).
+Concept Bottleneck Models (CBMs) [20] address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation).
 Predictions are then made as a linear combination of these concept activations.
 Initial CBMs required image–concept annotations to train the bottleneck layer that maps dense features to concepts.
-Modern CBMs (oikarinen2023labelfree; Yang2022LanguageIA; Panousis2023SparseLC; Rao2024DiscoverthenNameTC) overcome this limitation by leveraging the CLIP model (Radford2021LearningTV) as an annotation-free alternative.
-Since CLIP models map image and text into a shared embedding space, these approaches can query image features against a set of predefined textual concepts within that space and use cosine similarity scores to find a matching annotation.
+Modern CBMs [36, 60, 5] overcome this limitation by leveraging the CLIP model [43] or vision–language grounding models [48] to provide image-concept annotations.
+Since CLIP models map image and text into a shared embedding space, these approaches can query images against an entire pool of predefined textual concepts and use cosine similarity scores to find a matching annotation.
 These approaches are commonly referred to as label-free CBMs.
-Nevertheless, in many real-world scenarios a high‑performing, task‑specific legacy model already exists and typically outperforms zero‑shot CLIP models (zhang2024why).
-A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models?
+In many real-world settings, a high-performing, task-specific legacy model already exists and often achieves strong results on the target task [63].
+When CLIP is used to generate image-concept annotations, the resulting CBM is anchored to CLIP’s embedding space; the legacy model must be interpreted through CLIP’s notion of similarity, rather than through its own learned representation.
+In this sense, the approach is not fundamentally different from replacing the legacy model with a CLIP-based CBM.
+Furthermore, this dependence can transfer CLIP’s biases and behavior into the legacy model (e.g., importing the CLIP typographic bias [11] into a DINO model).
+A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models without the CLIP constraint?
 Retraining such a specialist on a large image–text corpus following the CLIP approach is impractical, in terms of computational cost and need of a huge amount of image-text data.
-Furthermore, obtaining ground-truth image-concept annotations is time consuming and labor intensive.
+Furthermore, obtaining ground-truth image-concept annotations through manual human-labor is time consuming and expensive.
 Finally, retraining this legacy model further alters its original decision-making process and distribution, which is typically not desired.
-As a result, CBMs remain confined to CLIP‑like models, leaving specialist vision models aside with no clear way to adapt them into the CBM framework.
-In this work, we lift this limitation by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP.
+In this work, we lift these limitations by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP.
 TextUnlock has four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier.
@@
 Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
-After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training but rather operating in a zero-shot manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
-In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP bottleneck (CLIP-free), neither annotated image-concept data (label-free) while also deriving the concept-to-class classifier in a zero-shot manner.
+After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training and operating in an unsupervised manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
+Figure 1: Overview of our proposed TextUnlock.
+(a) The process of training the MLP mapping between vision and text space.
+(b) The process of inference with the adapted visual classifier.
+The text encoder acts as weight generator for a linear classifier.
+Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
+In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP constraint (CLIP-free), neither on annotated image-concept data (label-free), and for the first time, also deriving the concept-to-class classifier in a fully unsupervised manner.
 (ii) We demonstrate the effectiveness of our method with 40 different architectures, along with extensive ablation studies and intervention results, and further show how our method can also be used to perform zero-shot image captioning.
-(iii) Our method sets new state-of-the-art results, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
+(iii) We set new state-of-the-art results on CBMs, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
 2 Related Work
-Concept Bottleneck Models (CBMs): Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) are a class of inherently interpretable models that map dense representations into human-understandable concepts, and then make predictions based on a linear combination of these interpretable units rather than opaque features.
-Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to perform the feature-to-concept mapping without requiring annotated image–concept data.
-This can be achieved either by using CLIP to provide ground-truth image–concept similarity scores that train the bottleneck layer (oikarinen2023labelfree), or by querying image features against a predefined set of concepts in the CLIP space (Yang2022LanguageIA; Panousis2023SparseLC) and using the resulting cosine similarities as concept activations.
-Several follow-up works (Rao2024DiscoverthenNameTC; Knab2024DCBMDV; Benou2025ShowAT; yuksekgonul2023posthoc) adopt this CLIP-based pipeline.
-However, these approaches remain limited to CLIP-based classifiers and cannot be applied to other visual models.
-Moreover, they require training a linear probe to map concept activations to class predictions.
-Our work addresses both limitations: we propose a fully CLIP-free CBM and demonstrate that our method can be readily used to derive the linear probe in a fully zero-shot manner.
+Concept Bottleneck Models (CBMs) are inherently interpretable models that predict by first mapping inputs to human-understandable concepts and then combining those concepts linearly to make the final decision, originally requiring concept annotations for each image [36].
+Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to provide image-concept annotations.
+CLIP is used to generate ground-truth image–concept similarity scores, which serve as supervision for training the bottleneck layer of a downstream classifier.
+Several further works [60, 5] adopted this pipeline.
+Another work uses vision-language grounding models [48] to obtain image-concept annotations.
+Another line of work [59, 39, ao2024DiscoverthenNameTC, 19] builds CBMs directly on top of CLIP, making them specifically tailored to CLIP-only models and thereby eliminating the need for a separate bottleneck training stage.
+In these methods, image features are queried against a predefined set of concepts in the CLIP embedding space, and the resulting cosine similarities are used directly as concept activations.
+As a result, all the above methods either (1) are limited to CLIP models, or (2) depend on CLIP or vision–language models to obtain the image-concept annotations.
+On the other hand, we propose a fully CLIP-free CBM; our method does not rely on any external vision-language model to generate the image-concept annotations, nor on human annotation, and can be applied to any legacy model without the CLIP constraint.
+Finally, all CBMs works to date require training a linear probe to map concept activations to class predictions.
+In contrast, we demonstrate that our method can be readily used to derive the linear probe in a fully unsupervised manner.
 Transforming Visual Features to Text: There exist works that try to decode visual features of models beyond CLIP.
-DeVIL (Dani2023DeViLDV) and LIMBER (merullo2023linearly) train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data.
+DeVIL [7] and LIMBER [34] train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data.
 Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process.
-ZS-A2T (Salewski2023ZeroshotTO) converts attention maps into natural language in a zero-shot manner using Large Language Models (LLMs), but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective).
-Text-to-Concept (T2C) (Moayeri2023TextToConceptB) also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
+Text-to-Concept (T2C) [35] also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder.
 Notably, all these methods rely on the CLIP approach and/or its supervision, and they alter the classifier’s distribution by entirely discarding its output class distribution.
@@
 3 Method
-We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design zero-shot concept bottleneck models, which we discuss in Section 3.2.
+We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design our proposed U-F2-CBM, which we discuss in Section 3.2.
+U stands for unsupervised; L2 for CLIP-free and Label-free (hence ‘double free’).
 3.1 TextUnlock
 A visual classifier assigns an image to a specific category from a predefined set of discrete class labels.
-For example, in ImageNet-1K (Deng2009ImageNetAL), this set contains $1,000$ class labels.
+For example, in ImageNet-1K [9], this set contains $1,000$ class labels.
 Originally, these discrete class labels correspond to class names in text format.
@@
 Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
-Figure 1: Overview of our proposed TextUnlock.
-(a) The process of training the MLP mapping between vision and text space with pseudocode given in Appendix Section A.
-(b) The process of inference with the adapted visual classifier.
-The text encoder acts as weight generator for a linear classifier.
-Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
 We propose to learn a lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$.
@@
 In other words, $S$ represents the classification logits of our approach.
+Figure 2: For any pretrained classifier, we first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in an unsupervised manner, which results in (c) our final U-F2-CBM.
+Note that the concept bank only needs to be encoded once.
 The most straightforward way to training the MLP is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution.
@@
 We instead propose to align $S$ to the original decision distribution $o$ through cross-entropy loss.
-For a single sample, the loss is given by
+For a single sample, the loss is
 $L=-\sum_{i=1}^{K}o_{i}\log\left(\frac{e^{s_{i}}}{\sum_{j=1}^{K}e^{s_{j}}}\right)\;.$ (1)
 This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but instead distill the distribution of the original model to its counterpart vision-language distribution.
-Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters.
+This loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters.
 Eq.
-1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose.
+1 shows that our approach does not require any labels for training, and steers the MLP to preserve the classifier’s distribution, since it is explicitly trained for that purpose.
 Note also that this loss function naturally encodes the classifier’s relationship across all classes.
-We provide an illustration of the TexUnlock approach in Figure 1 and a PyTorch-like pseudocode in Section A of the Appendix.
-It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section O in the Appendix for more information).
-After training, the projected visual features and the text encoder features lie in the same space.
-We can therefore query the visual features with any text by finding the alignment score between the embedded text and the projected visual features.
-In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the classification task formulated as argmax($\tilde{f}.U^{T}$), see Figure 1(b).
-3.2 Zero-Shot Concept Bottleneck Models
-Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed zero-shot CBMs.
-Note that at this stage, all model components (including the MLP) are frozen and no further training is performed.
-CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction.
-In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image.
-In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
-Concept Discovery.
-We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier.
-We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$.
-Following other works (Rao2024DiscoverthenNameTC), and without loss of generality, we use the $Z=20K$ most common words in English (mostcommongoogle) as our concept set.
-These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset.
-We ablate on other concept sets in Appendix Section I for the interested reader.
-To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set.
-Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”).
-In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
-Details of this procedure can be found in Section N of the Appendix.
-This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
-With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot.
-This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
-Figure 2: Building zero-shot CBMs for any pretrained classifier.
-We first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in a zero-shot manner, which results in (c) our final CBM.
-Note that the concept bank only needs to be encoded once.
-We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$.
-That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$.
-By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$.
-For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space.
-That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$.
-Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to.
-That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$.
-The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts.
-We provide an illustration in Figure 2(a).
-Concept-to-class Prediction.
-The classifier $W^{con}$ takes in concept activations and outputs a distribution $S_{cn}$ over classes.
-We build $W^{con}$ in a zero-shot manner, where here zero-shot indicates that no training is required to map concept activations to classes.
-Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space.
-Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name.
-Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$.
-That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$.
-Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts.
-This process is shown in Figure 2(b).
-In total, the output distribution $S_{cn}$ of the CBM is defined as:
-[…]
-From Eq.
-2 we make an interesting observation.
-Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
-The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other.
-Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$.
-Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model.
-Eq.
-2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts.
-This means our CBMs preserve the basic reasoning process of the original classifier.
-By this, we obtain zero-shot CBMs that discover concepts and build $W^{con}$ in a zero-shot manner for any classifier (Figure 2(c)).
-An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly.
-This makes our method highly flexible with respect to the chosen concept set.
-4 Experiments
-Model Top-1 Orig. $\Delta$ Model Top-1 Orig. $\Delta$
-ResNet50 75.80 76.13 -0.33 ResNeXt50-32x4dv2 80.79 80.88 -0.09
-[…]
-[…]
-ResNet101 77.19 77.37 -0.18 ViT-B/16 80.70 81.07 -0.37
-[…]
-[…]
-[…]
-[…]
-DenseNet169 75.46 75.60 -0.14 Swin-Base 83.22 83.58 -0.36
-EfficientNetv2-S 84.04 84.23 -0.19 Swinv2-Base 83.72 84.11 -0.39
-EfficientNetv2-M 84.95 85.11 -0.16 BeiT-B/16 84.54 85.06 -0.52
-[…]
-ConvNeXt-Small 83.42 83.62 -0.20 DINOv2-B 84.40 84.22 +0.18
-ConvNeXt-Base 83.88 84.06 -0.18 ConvNeXtV2-B 84.56 84.73 -0.17
-[…]
-[…]
-Table 1: TextUnlock-ed visual classifiers maintain classification performance.
-Comparison of our re-formulated classifiers for several models.
-Top-1 indicates our results of the new formulation, and Orig.
-denotes the original Top-1 accuracy.
-$\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
-We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock.
-We use the ImageNet-1K benchmark dataset due to the widespread publicly available visual classifiers trained and evaluated on it.
-In Section F of the Appendix, we also report results on other datasets including Places365 (zhou2017places), EuroSAT (helber2019eurosat) and DTD (Cimpoi2013DescribingTI) showing that our method is also applicable to domain-specific and fine-grained datasets.
-We apply TextUnlock on a diverse set of 40 visual classifiers.
-For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He2015DeepRL), Wide ResNets (Zagoruyko2016WRN), ResNeXts (Xie2016AggregatedRT), ShuffleNetv2 (Ma_2018_ECCV), EfficientNetv2 (Tan2021EfficientNetV2SM), Densely Connected Networks (DenseNets) (Huang2016DenselyCC), ConvNeXts (Liu2022ACF) and ConvNeXtv2 (Woo2023ConvNeXtVC).
-For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (dosovitskiy2021an), DINOv2 (oquab2024dinov), BeiT (bao2022beit), the hybrid Convolution-Vision Transformer CvT (Wu2021CvTIC), Swin Transformer (Liu2021SwinTH) and Swin Transformer v2 (Liu2021SwinTV).
-All models are pretrained on ImageNet-1K from the PyTorch (torchvision2016) and HuggingFace (wolf-etal-2020-transformers) libraries.
-Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K.
-Models with a subscript v2 are trained following the updated PyTorch training recipe (PytorchNewTrainRecepie).
-Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k.
-When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq.
-1.
-Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K.
-For the text encoder, we use the MiniLM Sentence Encoder (wang2020minilm) as it is powerful, fast and efficient.
-We provide ablation studies on other text encoders in Section D of the Appendix.
-Results are presented in Table 1.
-We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column.
-As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models.
-We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section C of the Appendix (see Section C.1 for the role and impact of the MLP and Section C.2 for its design hyperparameters).
-We also evaluate our transformed classifier robustness to prompt variations in Section H of the Appendix.
-For results on other models, we refer to Section L of the Appendix, and for implementation details, to Section J of the Appendix.
-[…]
-Effectivenss of Concept Interventions.
-Another common way to evaluate interpretability of CBMs is concept intervention.
-We report concept intervention results on our ZS-CBMs in Section E of the Appendix.
-In this way, we show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
-Supervised CBMs
-Method Model Top-1 Method Model Top-1
-LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
-LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
-CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
-DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
-DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
-DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
-Zero-Shot CBMs (Ours)
-Method Model Top-1 Method Model Top-1
-ZS-CBM ResNet50 73.9 ZS-CBM ViT-B/32 73.3
-[…]
-[…]
-[…]
-ZS-CBM WideResNet50 76.9 ZS-CBM Swinv2-Base 82.6
-[…]
-[…]
-ZS-CBM DenseNet121 69.9 ZS-CBM DINOv2-B 82.6
-[…]
-[…]
-ZS-CBM EfficientNetv2-M 83.9 ZS-CBM BeiT-L/16 86.2
-[…]
-[…]
-Table 2: Our zero-shot CBMs outperform CLIP-based counterparts.
-Accuracy of Supervised and Zero-Shot CBMs (ZS-CBMs, ours) on ImageNet validation set.
-Similar backbones are color-coded.
-Best within backbone family is underlined, overall best is bold.
-In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis.
-The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class.
-We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder.
-All examples use the LF-CBM concept set (oikarinen2023labelfree).
-By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab.
-Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature (Samek2019TowardsEA) regarding the “dumbbell” class.
-Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it.
-With our method, we can obtain a textual interpretation of the biases that the original classifier learns.
-We also provide qualitative examples of global class-wise concepts detected in Section P of the appendix.
-Figure 3: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the on the x-axis.
-5 Zero-Shot Image Captioning
-We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models.
-Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space.
-Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier.
-We adapt the method introduced in ZeroCap (Tewel2021ZeroCapZI) for our purpose.
-Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features.
-We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$.
-That is, $\tilde{f}=\text{MLP}(f)$.
-Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text.
-We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text.
-We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li2021PrefixTuningOC) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
-Specifically, we attach a set of learnable “virtual" tokens to $G$.
-Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number.
-$h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$.
-As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens.
-We perform this process for several iterations.
-As this process is not the core contribution of our work, we leave more details to Section M of the appendix, which also includes a detailed illustration in Figure 2.
-[…]
-Figure 4: Qualitative examples of zero-shot image captioning.
-Model B4 M R-L C S
-ZeroCap 2.6 11.5 14.6 5.5
-ConZIC 1.3 11.5 12.8 5.2
-Ours
-DenseNet161 1.50 10.2 20.4 15.8 6.3
-ResNet50 1.43 10.2 20.3 15.9 6.2
-WideResNet50 1.40 10.2 20.4 16.0 6.4
-[…]
-[…]
-[…]
-[…]
-DINOv2-Base 1.50 10.7 21.0 17.3 6.7
-[…]
-BeiT-L/16 1.50 10.6 20.9 17.6 6.9
-[…]
-[…]
-[…]
-Table 3: Zero-Shot Image Captioning Performance
-Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics.
-However, it is important to note that B4 and M are n-gram overlap-based metrics.
-They assume that the generated caption follows a specific structure and style.
-We verify this hypothesis by applying compositional image captioning (Kulkarni2011BabyTU; Lu2018NeuralBT) with in-context learning.
-Specifically, a set of image-grounded verbs and concepts (e.g., attributes, objects) are first detected using the concept discovery step described in Section 3.2, and they are fed along with six in-context examples from the COCO captioning set to a language model, capable of imitating the style and structure of the in-context examples, in order to compose them into a natural sounding sentence.
-This also demonstrates the direct application of our approach to compositional image captioning.
-The last row of Table 3 with the superscript com shows that the (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M compared to baseline methods.
-We provide more details and results on other models on this paradigm of compositional image captioning in Section K of the Appendix.
-We further show how we can use this approach to generate captions tailored to any domain (see Section K.1 in the Appendix).
-6 Conclusion
-We introduced a method for transforming any frozen visual classification model into a Concept Bottleneck Model (CBM).
-We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart.
-This method further enables a CLIP-free, label-free conversion of the frozen visual classification model into a CBM, and further allows us to derive zero-shot concept bottleneck concept-to-class classifiers (without training the linear classifier), Applied on 40 different models, our method outperforms supervised CLIP-based CBMs and achieves state-of-the-art results.
-TextUnlock further allows us to perform zero-shot image captioning on any visual classification model, also achieving state-of-the-art results and surpassing zero-shot captioning baselines based on CLIP.
-Finally, as with any research work, this study has its own limitations which are discussed in Section B of the Appendix.
-Appendix A Pseudocode of TextUnlock
-We provide a PyTorch-like pseudocode of TextUnlock in Listing 1.
+We provide an illustration of the TexUnlock approach in Figure 1, and a PyTorch-like pseudocode of it in Listing LABEL:lst:pseudo.
+It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section 14 in the Appendix for more information).
 # text_feats: textual features of class names from a frozen sentence encoder, shape (num_classes, text_dim)
@@
 Listing 1: PyTorch-like pseudocode for TextUnlock
-Appendix B Limitations
+After training, the projected visual features and the text encoder features lie in the same space.
+We can therefore query the visual features with any text by finding the alignment score between the embedded text and the projected visual features.
+In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the classification task formulated as argmax($\tilde{f}.U^{T}$), see Figure 1(b).
+3.2 U-F2-CBMs
+Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed U-F2-CBMs.
+Note that at this stage, all model components (including the MLP) are frozen and no further training is performed.
+A CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction.
+In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image.
+In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
+Concept Discovery: We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which represent the classification weights of the newly formulated classifier.
+In CBMs, we are given a large set of textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$.
+Following other works [44] and without loss of generality, we use the $Z=20K$ most common words in English [12] as our concept set.
+These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset.
+We ablate on other concept sets in Appendix Section 7 for the interested reader.
+To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set.
+Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”).
+In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
+Details of this procedure can be found in Section 13 of the Appendix.
+This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
+With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot.
+This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
+We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$.
+That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$.
+By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$.
+For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space.
+That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$.
+Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to.
+That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$.
+The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts.
+Concepts that the model identifies in an image will produce a high activation score.
+We provide an illustration in Figure 2(a).
+Model Top-1 Orig. $\Delta$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: TextUnlock-ed visual classifiers for CNNs and Transformers maintain classification performance.
+Comparison of our re-formulated classifiers for several models (remaining models in Appendix).
+Top-1 indicates our results of the new formulation, and Orig.
+denotes the original Top-1 accuracy.
+$\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
+Concept-to-Class Prediction: The classifier $W^{con}$ takes the concept activations produced from the concept discovery stage, and outputs a distribution $S_{cn}$ over classes.
+We build $W^{con}$ in an unsupervised manner; here unsupervised indicates that no training is required to map concept activations to classes.
+Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space.
+Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name.
+Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$.
+That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$.
+Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts.
+This process is shown in Figure 2(b).
+While the operation is performed entirely in the text feature space, the text encoder serves as the classifier weight generator.
+Because the MLP learns to map image representations to the text-encoder space, it is effectively mapping images into the classifier weight space.
+In total, the output distribution $S_{cn}$ of the CBM is obtained by feeding the identified concept activations in the image obtained from the concept discovery stage, to $W^{con}$.
+That is,
+[…]
+From Eq.
+2 we make an interesting observation.
+Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$).
+The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other.
+Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$.
+Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model.
+Eq.
+2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts.
+This means our CBMs preserve the basic reasoning process of the original classifier.
+By this, we obtain CBMs that discover concepts and build $W^{con}$ in an unsupervised manner for any classifier (Figure 2(c)).
+An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly.
+This makes our method highly flexible to the concept set.
+Supervised CBMs
+Method Model Top-1 Method Model Top-1
+LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
+LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
+CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
+DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
+DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
+DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
+Unsupervised, Label-Free, CLIP-Free CBMs (U-F2-CBM)
+Method Model Top-1 Method Model Top-1
+U-F2-CBM ResNet50 73.9 U-F2-CBM ViT-B/32 73.3
+[…]
+[…]
+[…]
+U-F2-CBM WideResNet50 76.9 U-F2-CBM Swinv2-Base 82.6
+[…]
+[…]
+U-F2-CBM DenseNet121 69.9 U-F2-CBM DINOv2-B 82.6
+[…]
+[…]
+U-F2-CBM EfficientNetv2-M 83.9 U-F2-CBM BeiT-L/16 86.2
+[…]
+[…]
+Table 2: Our U-F2-CBMs outperform CLIP-based counterparts.
+Accuracy of Supervised and U-F2-CBMs (ours) on ImageNet validation set.
+Similar backbones are color-coded.
+Best within backbone family is underlined, overall best is bold.
+4 Experiments
+We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock.
+We use the most challenging CBM benchmark of ImageNet-1K dataset.
+There is also widespread publicly available visual classifiers trained and evaluated on ImageNet.
+In Section 5 of the Appendix, we also report results on other datasets including Places365 [64], EuroSAT [14] and DTD [6] showing that our method is also applicable to domain-specific, fine-grained datasets as well as datasets with a small number of classes.
+We apply TextUnlock on a diverse set of 40 visual classifiers.
+For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [13], Wide ResNets [61], ResNeXts [58], ShuffleNetv2 [31], EfficientNetv2 [50], Densely Connected Networks (DenseNets) [17], ConvNeXts [28] and ConvNeXtv2 [56].
+For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [10], DINOv2 [38], BeiT [4], the hybrid Convolution-Vision Transformer CvT [57], Swin Transformer [27] and Swin Transformer v2 [26].
+All models are pretrained on ImageNet-1K from the PyTorch [32] and HuggingFace [55] libraries.
+Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K.
+Models with a subscript v2 are trained following the updated PyTorch training recipe [53].
+BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k.
+When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq.
+1.
+Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K.
+For the text encoder, we use the MiniLM Sentence Encoder [54] as it is fast and efficient.
+We provide ablation studies on other text encoders in Section 2 of the Appendix.
+Results for 17 classifiers are presented in Table 1, with the remaining 23 classifiers reported in Section 11 of the Appendix.
+We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column.
+As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models.
+We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section 3 of the Appendix (see Section 3.1 for the role and impact of the MLP and Section 3.2 for its design hyperparameters).
+We also evaluate our transformed classifier robustness to prompt variations in Section 6 of the Appendix.
+For implementation details, we refer to Section 8 of the Appendix.
+[…]
+Other Datasets:
+We experiment with other datasets and show that our method is applicable even to domain-specific, fine-grained datasets as well as datasets with a small number of classes.
+We report CBM results on Places365 (domain-specific to scenes, 365 classes), DTD (domain-specific to texture and fine-grained, 47 classes), and EuroSAT (domain-specific to satellite images, 10 classes).
+When a baseline method is also reported on that dataset, we include it.
+Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline.
+All baselines use the same concept set.
+Note also that all baselines train a supervised linear classifier, while our method derives the classifier in an unsupervised manner.
+Results are shown in Table 3.
+For Places 365, we can see that U-F2-CBM for DenseNet161 classifier (using the transformed ImageNet-only trained classifier with TextUnlock) outperforms supervised CLIP-based ResNet and ViTs CBM methods.
+The same applies for EuroSAT and DTD with our U-F2-CBM.
+Therefore, the experiments show that our method scales to those scenarios as well.
+Dataset Method Model Acc (%)
+Places365 DCLIP CLIP-ResNet50 37.90
+DCLIP CLIP-ViT-B/16 40.30
+LF-CBM CLIP-ResNet50 49.00
+LF-CBM CLIP-ViT-B/16 50.60
+CDM CLIP-ResNet50 52.70
+CDM CLIP-ViT-B/16 52.60
+Ours ResNet50 51.57
+Ours DenseNet161 53.42
+EuroSAT Baseline CLIP-ResNet50 86.27
+Baseline CLIP-ViT-B/16 88.57
+Ours ViT-B/16 93.65
+Ours WideResNet101 94.12
+Ours ResNet50 94.22
+DTD Baseline CLIP-ResNet50 57.77
+Baseline CLIP-ViT-B/16 61.86
+Ours WideResNet101 66.97
+Ours ViT-B/16 68.46
+Ours ResNet50 68.88
+Table 3: CBM results on Places365, EuroSAT, and DTD datasets.
+Figure 3: Qualitative examples of our U-F2-CBM. We show the top-detected concepts, each with their corresponding importance score.
+Concept Interventions:
+Another common way to evaluate interpretability of CBMs is concept intervention.
+This follows metrics from the explainability literature [41].
+In this way, we can show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
+As these metrics serve as complementary evaluations, we refer readers to Section 4 of the Appendix for the results.
+In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis.
+The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class.
+We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder.
+All examples use the LF-CBM concept set [36].
+By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab.
+Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [47] regarding the “dumbbell” class.
+Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it.
+With our method, we can obtain a textual interpretation of the biases that the original classifier learns.
+We also provide qualitative examples of global class-wise concepts detected in Section 15 of the appendix.
+5 Zero-Shot Image Captioning
+We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models.
+Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space.
+Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier.
+We adapt the method introduced in ZeroCap [51] for our purpose.
+Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features.
+We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$.
+That is, $\tilde{f}=\text{MLP}(f)$.
+Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text.
+We utilize an off-the-shelf pretrained language decoder model, denoted as $G$, to generate open-ended text.
+We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [22] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$.
+More details about ZeroCap is in Section 12 of the appendix.
+[…]
+Model B4 M R-L C S
+ZeroCap 2.6 11.5 14.6 5.5
+ConZIC 1.3 11.5 12.8 5.2
+Ours
+ResNet50 1.43 10.2 20.3 15.9 6.2
+[…]
+[…]
+DINOv2-Base 1.50 10.7 21.0 17.3 6.7
+BeiT-L/16 1.50 10.6 20.9 17.6 6.9
+[…]
+[…]
+[…]
+Table 4: Zero-Shot Image Captioning Performance
+6 Conclusion
+We introduced a method for transforming any frozen visual classification model into a CBM.
+We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart.
+This allows us to then produce an unsupervised, label-free and CLIP-free (U-F2) CBM, which outperforms supervised CLIP-based CBMs across 40 models.
+We also showed zero-shot image captioning as an additional application.
+Finally, as with any research work, this study has its own limitations, discussed in Section 1 of the Appendix.
+Supplementary Material
+1 Limitations
 As with any research work, our study has its own limitations that should be transparent and acknowledged.
-In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs.
+In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs, due to polysemy.
 This issue is closely tied to the choice of concept set used.
 When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts.
-Figure 1 illustrates such cases.
-In the first example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric).
+For example, the bird “drake" is linked to artist-related concepts (Rihanna, Robbie, lyric).
 This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake".
 In fact, a google search with the word “drake” yields directly the artist rather than the bird.
-In the second example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself.
+As another example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself.
 We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms).
-For each example, we also report the total logit score of the prediction.
 However, it is worth noting the following:
@@
     This issue also appears in CLIP-based CBMs and hence not unique to our approach.
-Figure 1: Limitations of our method in wrong semantic concept association
-Appendix C Ablation Studies of the MLP
-C.1 Role and Impact of the MLP
+2 Ablation Studies on the Text Encoder
+In Table 1, we present ablation studies using other text encoders from the Sentence-BERT library [45] with the ResNet50 visual classifier.
+We observe that the choice of the text encoder has minimal effect on the performance.
+This is because even lower-performing text encoders are capable of understanding class names.
+Text Encoder Top-1 (%)
+DistilRoberta 75.73
+MPNet-Base 75.78
+MPNet-Base-MultiQA 75.76
+MiniLM 75.80
+Table 1: Ablation studies on other text encoders
+3 Ablation Studies of the MLP
+3.1 Role and Impact of the MLP
 We first evaluate the role and impact of the MLP to verify its contribution and that it learns meaningful transformations, and does not collapse into a trivial function.
@@
 Table 2: Ablation studies of the MLP.
-C.2 MLP Design
+3.2 MLP Design
 Next, we perform an ablation study over the number of layers and output dimension scaling factor (dim_out_factor) of the MLP.
@@
 Table 3: Performance comparison of the MLP for different layer and dimension configurations.
-Appendix D Ablation Studies on the Text Encoder
-In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library (reimers-2019-sentence-bert) with the ResNet50 visual classifier.
-We observe that the choice of the text encoder has minimal effect on the performance.
-This is because even lower-performing text encoders are capable of understanding class names.
-Text Encoder Top-1 (%)
-DistilRoberta 75.73
-MPNet-Base 75.78
-MPNet-Base-MultiQA 75.76
-MiniLM 75.80
-Table 3: Ablation studies on other text encoders
-Appendix E Effectiveness of Concept Interventions
-As a supplementary experiment, we also report concept intervention results on our ZS-CBMs.
+4 Concept Interventions and other Explainability Metrics
+4.1 Concept Interventions
+We report concept intervention results on our ZS-CBMs.
 Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
-The Waterbirds-100 dataset (Sagawa2020Distributionally) is a standard dataset used in previous works (Rao2024DiscoverthenNameTC) to conduct CBM intervention experiments.
+The Waterbirds-100 dataset [46] is a standard dataset used in previous works [44] to conduct CBM intervention experiments.
 It is a binary classification dataset of two classes: waterbirds and landbirds.
@@
 By building a CBM, we can correct this bias by intervening in concepts in the CB layer.
-We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set.
-Our curated dataset includes 140 validation images (70 for each class).
-We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from (Rao2024DiscoverthenNameTC) which includes a collection of bird-related concepts and a collection of land-related concepts.
+We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set, including 140 validation images (70 for each class).
+We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from [44] which includes a collection of bird-related concepts and a collection of land-related concepts.
 The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification.
-To correct this, we intervene in the concepts in the CB layer, following the setup from (Rao2024DiscoverthenNameTC).
+To correct this, we intervene in the concepts in the CB layer, following the setup from [44].
 Intervention R (Int.
 R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop.
+The larger the drop, the better.
 Intervention K (Int.
 K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase.
+More increase is better.
 Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
@@
 A drop in accuracy indicates that these concepts are important for prediction.
-We select a subset of 10 classes from ImageNet following (Howard_Imagenette_2019).
-We use the original ImageNet validation images for those classes, rather than the validation set from (Howard_Imagenette_2019), because the latter contains original ImageNet training examples that our model has already seen.
+We select a subset of 10 classes from ImageNet following [16].
+We use the original ImageNet validation images for those classes, rather than the validation set from [16], because the latter contains original ImageNet training examples that our model has already seen.
 We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts.
-The classes and concepts we used are provided in Section Q of the supplementary material.
+The classes and concepts we used are provided in Section 16 of the supplementary material.
 We then measure CBM accuracy before and after intervention.
@@
 Table 5: CBM Interventions in a multi-class setup
-Appendix F Transformed Classifier Results on Other Datasets
+4.2 Concept Discovery Accuracy
+We use the discrete-based Mutual Information (MI) metric [^2] https://github.com/fawazsammani/clip-interpret-mutual-knowledge to calculate the MI between concepts detected by our method, and those detected by CLIP.
+Specifically, given the concept bank, we use the CLIP Vision Encoder to encode an image, and the CLIP Text Encoder to encode the concept bank.
+We record the set of top-K detected concepts most similar to the image.
+We then encode the same image with our TextUnlock-ed visual classifier (e.g., DINOv2)(including the MLP) and encode the same concept bank with our Text Encoder $T$.
+We record the set of top-K detected concepts most similar to the image.
+We then calculate the MI between the detected concepts from CLIP (CLIP ViT-B/16 model) and our method.
+Results are shown in Table 6 and demonstrate high mutual information which indicates that our method successfully detects key concepts in the dataset by only learning from the limited vocabulary of class labels.
+Model NMI
+DINOv2 0.681
+EfficientNetv2-M 0.704
+ViT-B/16v2 0.711
+Table 6: Normalized Mutual Information (NMI) with CLIP ViT-B/16 for different models. NMI ranges from 0 (no mutual information) to 1 (perfect agreement).
+5 Transformed Classifier Results on Other Datasets
 We conduct extra experiments on 3 datasets.
@@
 Table 7: Results of our transformed classifier on Places365, EuroSAT, and DTD datasets
-Appendix G CBM Results on Other Datasets
-We report CBM results on additional datasets: Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images).
-When a baseline method is also reported on that dataset, we include it as a baseline.
-Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline.
-All baselines use the same concept set.
-Results are shown in Table 7.
-For Places 365, we can see that the transformed DenseNet161 classifier trained only on ImageNet, outperforms supervised CLIP-based ResNet and ViTs CBM methods in a zero-shot manner.
-The same applies for EuroSAT and DTD with our transformed Image-Net only trained classifiers.
-Therefore, the experiments show that our method scales to domain-specific and fine-grained datasets.
-Dataset Method Model Acc (%)
-Places365 Supervised
-DCLIP CLIP-ResNet50 37.90
-DCLIP CLIP-ViT-B/16 40.30
-LF-CBM CLIP-ResNet50 49.00
-LF-CBM CLIP-ViT-B/16 50.60
-CDM CLIP-ResNet50 52.70
-CDM CLIP-ViT-B/16 52.60
-Zero-Shot
-Ours ResNet50 51.57
-Ours DenseNet161 53.42
-EuroSAT Supervised
-Baseline CLIP-ResNet50 86.27
-Baseline CLIP-ViT-B/16 88.57
-Zero-Shot
-Ours ResNet50 94.22
-Ours WideResNet101 94.12
-Ours ViT-B/16 93.65
-DTD Supervised
-Baseline CLIP-ResNet50 57.77
-Baseline CLIP-ViT-B/16 61.86
-Zero-Shot
-Ours ResNet50 68.88
-Ours WideResNet101 66.97
-Ours ViT-B/16 68.46
-Table 7: CBM results on Places365, EuroSAT, and DTD datasets
-Appendix H Prompt Variations
+6 Prompt Variations
 We evaluate the robustness of our transformed classifier to variations in text prompts.
@@
 Table 8: Effect of different text prompts on Top-1 accuracy using ViT-B/16.
-Appendix I Other Concept Sets
+7 Other Concept Sets
 Early works such as label-free CBMs use LLMs to automatically generate a set of concepts for a target class.
-In recent works such as DN-CBM (Rao2024DiscoverthenNameTC) and DCBM (Knab2024DCBMDV), it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations.
+In recent works such as DN-CBM [44] and DCBM [19], it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations.
 In Table 9, we compare general pre-defined and LLM-generated concept sets.
@@
 Table 9: Comparison of Predefined and LLM-generated concept sets on CBMs
-Appendix J Implementation Details
-For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (reimers-2019-sentence-bert).
+8 Implementation Details
+For the text encoder, we use the all-MiniLM-L12-v1333https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library [45].
 This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs.
 It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
-Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba2016LayerN), a GELU activation function (Hendrycks2016GaussianEL) and Dropout (Srivastava2014DropoutAS) with a drop probability of 0.5.
+Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization [2], a GELU activation function [15] and Dropout [49] with a drop probability of 0.5.
 The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function.
 The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder).
-We train the MLP projector with a batch size of 256 using the ADAM optimizer (KingBa15) with a learning rate of 1e-4 that decays using a cosine schedule (loshchilov2017sgdr) over the total number of epochs.
+We train the MLP projector with a batch size of 256 using the ADAM optimizer [18] with a learning rate of 1e-4 that decays using a cosine schedule [29] over the total number of epochs.
 We follow the original image sizes that the classifier was trained on.
@@
 Each model is trained on a single NVIDIA GeForce RTX 2080 Ti GPU.
-Appendix K Compositional Image Captioning
-Compositional Image Captioning is an old image captioning paradigm termed (Kulkarni2011BabyTU), later revived with deep learning methods (Lu2018NeuralBT).
+9 Compositional Image Captioning
+Compositional Image Captioning is an old image captioning paradigm termed [21], later revived with deep learning methods [30].
 In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence.
@@
 For the concepts, we use the same concept set as in Section 3.2.
-We also add a list of the most common verbs (commonverbs) in English to the pool.
+We also add a list of the most common verbs [8] in English to the pool.
 This allows us to cover all possible words and interactions.
-We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 9 examples).
-This allows us to generate sentences adhering to a specific style and structure.
-For this experiment, we used GPT4o-mini (gpt4omini) as our LLM, as it is fast and cost-efficient.
+We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 6 examples).
+This allows us to generate sentences adhering to a specific style and structure of the in-context examples.
+For this experiment, we used GPT4o-mini [37] as our LLM, as it is fast and cost-efficient.
 In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importance.
@@
 While the results on CiDEr and SPICE are incremental compared to the results in Table 4, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis that the low scores of B4 and M were attributed to the specific caption style and structure of the respective dataset (COCO).
-K.1 Domain-Specific Compositional Captioning
+9.1 Domain-Specific Compositional Captioning
 Since the LLM we used for compositional image captioning is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains.
@@
 Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset.
-We use the ImageNet-specific concept set from (oikarinen2023labelfree) and report zero-shot captioning performance in Table 11.
+We use the ImageNet-specific concept set from [36] and report zero-shot captioning performance in Table 11.
 As shown, we observe a decrease in all metrics.
@@
 Table 11: Composition Captioning Performance using the ImageNet-specific LF-CBM concept set
-Appendix L Performance on Additional Models
+10 Qualitative examples of zero-shot image captioning.
+We present qualitative examples of zero-shot image captioning from different classifiers in Figure 1.
+We choose BeiT-L/16, ConvNeXtv2, and ViT-B/16.
+This allows us to see how different classifiers “see" the image.
+From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables.
+In contrast, ViT-B/16 focuses exclusively on the dog and its characteristics.
+Figure 1: Qualitative examples of zero-shot image captioning.
+11 Performance on Additional Models
 We report performance on additional models that were not included in the main manuscript in Table 12.
 Model Top-1 Orig. $\Delta$
+[…]
+BeiT-B/16 84.54 85.06 -0.52
+WideResnet50 78.35 78.47 -0.12
+[…]
+EfficientNetv2-S 84.04 84.23 -0.19
+ConvNeXt-Small 83.42 83.62 -0.20
+ResNeXt101-32x8d 79.10 79.31 -0.21
+[…]
+[…]
+DenseNet169 75.46 75.60 -0.14
 ConvNeXt-Tiny 82.19 82.52 -0.33
+[…]
+[…]
+ResNeXt50-32x4d 77.44 77.62 -0.18
+ConvNeXt-Base 83.88 84.06 -0.18
+ResNeXt50-32x4dv2 80.79 80.88 -0.09
 ViT-B/32 75.40 75.91 -0.51
@@
 Swinv2-Small 83.32 83.71 -0.39
-[…]
 Table 12: Performance of our reformulated classifiers for additional models
-Appendix M Process of Zero-Shot Image Captioning
+12 Process of Zero-Shot Image Captioning
 We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$.
 A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$.
-To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li2021PrefixTuningOC), which prepends learnable tokens in the embedding space.
+To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning [22], which prepends learnable tokens in the embedding space.
 We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly.
-Our method builds upon the work of (Tewel2021ZeroCapZI).
+Our method builds upon the work of [51].
 A high-level overview of this process is illustrated in Figure 2.
@@
 The final output is chosen as the sentence with the highest similarity to the visual features $\tilde{f}$.
-We also add the fluency loss from (Tewel2021ZeroCapZI) as well as other token processing operations.
-We refer readers to (Tewel2021ZeroCapZI) for more information.
-We use the small GPT-2 of 124M parameters as $G$.
+We also add the fluency loss from [51] as well as other token processing operations.
+We refer readers to [51] for more information.
+We use the smallest GPT-2 of 124M parameters as $G$.
 We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
@@
 We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features.
-Appendix N Concept Filtering
+13 Concept Filtering
 One of our concept filtering procedures requires us to know the terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name.
@@
 Answer: fish,animal,cyprinid,carp,vertebrate.
-Appendix O Using only class names
+14 Using only class names
 We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP.
-In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin1998WordNetAE) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt2022WhatDA) or VCVD(Menon2022VisualCV).
+In practice, we can go beyond class names by using resources like a class hierarchy from WordNet [24] (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL [42] or VCVD[33].
 However, this approach would be considered as “cheating.
@@
 We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
-Appendix P Global Class-Wise Qualitative Examples
+15 Global Class-Wise Qualitative Examples
 In Figure 3, we present a probability distribution of global class-wise concepts.
@@
 These concepts are “harpoon" and “lobster hammer", both which are distinctive to the head of the hammerhead shark and drive its prediction.
+Note that for this experiment, we do not apply the concept filtering procedure.
 Figure 3: Global class-wise interpretability analysis with our Concept Bottleneck Model.
 We highlight in yellow the top concepts in "hammerhead shark" that are not present in "tiger shark", and therefore distinctive to "hammerhead shark".
-Appendix Q Multi-class CBM Intervention Classes and Concepts
+16 Multi-class CBM Intervention Classes and Concepts
 Class Concepts